### v1

# Does it matter? A more careful treatment of density fluctuations in 21-cm simulations

Abstract
The cosmological 21-cm signal is sourced from hyperfine transitions in neutral hydrogen atoms. Yet, although hydrogen atoms belong to the sub-class of baryonic matter, it is customary to simplify their treatment and analyze the 21-cm signal as if all the matter in the Universe was in the form of collisionless cold dark matter (CDM). This is usually done by evolving the density field via a scale-independent growth factor (SIGF). In this work, we separate the baryons from CDM and evolve the two species with a proper scale-dependent growth factor (SDGF). By incorporating the SDGF in the 21cmFirstCLASS code, we demonstrate the effect that baryons and CDM have on the 21-cm signal at the linear dark ages epoch and the subsequent non-linear epochs of cosmic dawn and reionization. Our analysis shows that the baryonic nature of hydrogen cannot be ignored during the dark ages, and that non-linear effects in density-field evolution must be accounted for after stars have formed. Furthermore, we discuss how the 21-cm signal is modified at lower redshifts, where ground-based 21-cm interferometers are mostly sensitive, due to the choice of working with either the “linear” or “non-linear” matter over-density (that is, the over-density as computed from linear perturbation theory, and non-linear perturbation theory, respectively) in the extended Press-Schechter formalism. Our code is publicly available at https://github.com/jordanflitter/21cmFirstCLASS.
I Introduction
Between recombination and reionization, the Universe was filled with neutral hydrogen atoms. Hyperfine energy de-excitations in these atoms, following interactions with other particles, led to the emission of 21-cm wavelength photons. This is the 21-cm signal in a nutshell Madau et al. (1997); Barkana and Loeb (2001); Loeb and Zaldarriaga (2004); Bharadwaj and Ali (2005); Furlanetto et al. (2006); Pritchard and Loeb (2012); Bera et al. (2023); Shaw et al. (2023). This exciting signal has the potential to expand our knowledge on the formation mechanism of the first galaxies Barkana and Loeb (2005a); Furlanetto (2006); Park et al. (2019); Muñoz et al. (2020); Park et al. (2020); Greig et al. (2020a); Thyagarajan et al. (2020); Qin et al. (2020, 2021a); Reis et al. (2022a); Ma et al. (2023), the thermal history of the intergalactic medium (IGM) Chen and Miralda-Escude (2004); Mesinger et al. (2014); Ghara and Mellema (2020); Greig et al. (2021); Ghara et al. (2021); Maity and Choudhury (2022); Abdurashidova et al. (2022, 2023); Lazare et al. (2024a), the epoch of reionization Lidz et al. (2009); Wiersma et al. (2013); Jensen et al. (2013); Meerburg et al. (2013); Greig and Mesinger (2015, 2017); Greig et al. (2020b); Ghara et al. (2020); Watkinson et al. (2022); Rahimi et al. (2021); Qin et al. (2021b); Saxena et al. (2023); Giri et al. (2024); Ghara et al. (2024), physics beyond the standard model Tashiro and Sugiyama (2006); Schleicher et al. (2009); Shiraishi et al. (2014); Chluba et al. (2015); Villaescusa-Navarro et al. (2015); Oyama et al. (2016); Pal and Guha Sarkar (2016); Kovetz et al. (2019); Caputo et al. (2021); Natwariya and Nayak (2022); Kunze (2019); Minoda et al. (2022); Shmueli et al. (2023); Kunze (2023); Cruz et al. (2024a); Libanore et al. (2024); Bhaumik et al. (2024); Lee and Hotinli (2024); Dhuria and Teli (2024); Plombat et al. (2024); Adi et al. (2024), and possibly shed some light on the nature of dark matter Tashiro et al. (2014); Evoli et al. (2014); Muñoz et al. (2015a); Lopez-Honorez et al. (2016); D’Amico et al. (2018); Cheung et al. (2019); Mitridate and Podo (2018); Lidz and Hui (2018); Nebrin et al. (2019); Fialkov et al. (2018); Barkana (2018); Clark et al. (2018); Kovetz et al. (2018); Berlin et al. (2018); Barkana et al. (2018); Slatyer and Wu (2018); Muñoz and Loeb (2018); Liu and Slatyer (2018); Muñoz et al. (2018); Creque-Sarbinowski et al. (2019); Liu et al. (2019); Hotinli et al. (2022); Jones et al. (2021); Halder et al. (2021); Sarkar et al. (2022); Barkana et al. (2023); Giri and Schneider (2022); Flitter and Kovetz (2022); Lazare et al. (2024b); Driskell et al. (2022); Dey et al. (2023); Short et al. (2022); Mondal et al. (2023); Facchinetti et al. (2024); Qin et al. (2024); Sun et al. (2023); Vanzan et al. (2024); Cang et al. (2023); Shao et al. (2023); Katz et al. (2024); Lopez-Honorez (2024).
For these reasons, worldwide efforts are being made to measure the 21-cm signal by various collaborations and experiments. Some of these experiments are designed to measure the “global” (sky-averaged) 21-cm signal, like the Experiment to Detect the Global reionization Signature (EDGES) Monsalve et al. (2019), Shaped Antenna measurement of the background RAdio Spectrum (SARAS) Nambissan T. et al. (2021), Large-Aperture Experiment to Detect the Dark Ages (LEDA) Price et al. (2018), the Radio Experiment for the Analysis of Cosmic Hydrogen (REACH) de Lera Acedo et al. (2022) and Probing Radio Intensity at high-Z from Marion (PRizM) Philip et al. (2019). Other experiments, like the Hydrogen Epoch of Reionization Array (HERA) DeBoer et al. (2017), Low Frequency Array (LOFAR) van Haarlem et al. (2013), Murchison Widefield Array (MWA) Tingay et al. (2013), the Giant Metrewave Radio Telescope (GMRT) Swarup et al. (1991), the Precision Array for Probing the Epoch of Reionization (PAPER) Kolopanis et al. (2019) and Square Kilometre Array (SKA) Braun et al. (2015) have been dedicated to measuring the spatial fluctuations in the signal using ground based interferometers. All the aforementioned experiments are limited to detecting the signal at redshifts $z\lesssim 30$ (namely, high frequencies), as the Earth ionosphere distorts incoming low-frequency radiation that is sourced at very high redshifts. Several interesting proposals are considered nowadays to overcome this obstacle by deploying radio receivers on the far side of the Moon, or in orbit around it or the Earth Jester and Falcke (2009); Burns et al. (2021a); Furlanetto et al. (2019); Silk (2021); Burns et al. (2021b, a); Goel et al. (2022); Shi et al. (2022); Bertone et al. (2023); Fialkov et al. (2023).
The growing interest in the 21-cm signal and its analysis motivated the development of a plethora of dedicated codes, designed to estimate the signal and its spatial fluctuations as a function of redshift. After recombination, prior to star formation, the fluctuations in the signal are tiny and linear perturbation theory can be applied in that regime. For example, CAMB Lewis and Challinor (2007) and PowerSpectrum21cm Muñoz et al. (2015b) exploit the linearity of the perturbations in this “dark ages” epoch. After cosmic dawn, once galaxies and stars have formed, the fluctuations in the signal are amplified considerably, and the validity of linear perturbation theory breaks. Different techniques have been adapted in the literature to simulate the signal in the non-linear regime of cosmic dawn and reionization. This includes full radiative transfer simulations (CodA Ocvirk et al. (2016, 2020); Lewis et al. (2022), 21SSD Semelin et al. (2017), THESAN Kannan et al. (2022)), one-dimensional radiative transfer methods (BEoRN Schaeffer et al. (2023), Bears Thomas et al. (2009), Grizzly Ghara et al. (2018)), hydrodynamic-radiative-transfer frameworks (Licorice Semelin et al. (2007); Meriot and Semelin (2024)), and ray-tracing algorithms (C2-ray Mellema et al. (2006); Hirling et al. (2023), CRASH Maselli et al. (2003)).
Figure 1: The global 21-cm signal under different assumptions. Green lines are obtained using 21cmFirstCLASS and separate the baryons and dark matter evolution; other lines assume $\delta_{b}\equiv\delta_{c}$. In the blue curve, we ran 21cmFAST under the default settings of non-linear evolution; the orange curve instead is obtained setting the density field to evolve linearly, similarly as in Zeus21 (the two codes well agree in these settings). Solid (dashed) lines correspond to codes that adopt the non-linear (linear) matter density field in their calculations. In both green curves of 21cmFirstCLASS, the baryon density field is evolved non-linearly.
Figure 2: Comparison of the 21-cm power spectrum on different scales, under the same conditions as in Fig. 1. Here, we show with error bars the expected noise from HERA in its design sensitivity at low redshifts ($z<10$) assuming the moderate foreground scenario, as evaluated by the 21cmSense code Pober et al. (2013, 2014) (we used the same 21cmSense settings as detailed in Ref. Flitter and Kovetz (2024a)). We also show a separate error bar at $z=14$, corresponding to a sensitivity estimate for SKA Meriot and Semelin (2024).
Semi-numerical codes, whose runtime is shorter by orders of magnitude, are also found in the literature. Among these, 21cmFAST Mesinger et al. (2011); Muñoz et al. (2022) is perhaps the most popular one, having also been used to analyse HERA Phase I data Abdurashidova et al. (2022, 2023), but see also simfast21 Santos et al. (2010), CIFOG Hutter (2018) and 21cmSPACE Visbal et al. (2012); Fialkov et al. (2013, 2014a, 2014b); Fialkov and Barkana (2014); Fialkov et al. (2015); Cohen et al. (2016); Fialkov et al. (2017); Fialkov and Barkana (2019); Reis et al. (2020, 2021, 2022b); Magg et al. (2022); Gessey-Jones et al. (2022); Sikder et al. (2023); Gessey-Jones et al. (2023); Pochinda et al. (2024), though the latter is not publicly available as of yet. Recently, using output from 21cmFAST, and assuming linear evolution of the density field, a fully analytic code named Zeus21 Muñoz (2023); Cruz et al. (2024b) was calibrated at the cosmic dawn era, allowing the code to compute the 21-cm power spectrum in mere seconds. [^1] ARES Mirocha (2014); Mirocha et al. (2017) is another speedy code, but unlike Zeus21 and HMreio, it only computes the global 21-cm signal. Another analytic code, named HMreio Schneider et al. (2023), also calibrated to 21cmFAST, focuses on the epoch of reionization. All these codes use a single quantity $\delta$ to denote over-density, and thus implicitly assume $\delta_{b}=\delta_{c}=\delta_{m}$, namely that baryons (“$b$”) and cold dark matter (CDM, “$c$”) behave alike (“$m$” stands for all of matter).
In this paper, we relax this assumption and explore the cosmological implications of distinguishing $\delta_{b}$ from $\delta_{c}$ in the context of the 21-cm signal. To do so, we rely on the newly developed tool 21cmFirstCLASS Flitter and Kovetz (2024a, b). This code interfaces CLASS Blas et al. (2011) and 21cmFAST, allowing the simulation to begin from recombination with initial conditions that are consistent with the simulated cosmology. We find that at the dark ages the approximation $\delta_{b}\approx\delta_{c}$ completely fails, and $\delta_{b}$ must be evolved in a scale-dependent manner in that epoch. We also demonstrate, using scale-dependent Lagrangian perturbation theory (LPT), the need to go beyond linear perturbation theory for the baryon over-density field at low redshifts, after cosmic dawn was set. Finally, we present how the signal is modified when considering either the “linear” or “non-linear” matter over-density $\delta_{m}$ in the context of the extended Press-Schechter formalism, and we discuss the role of the relative velocity between baryons and dark matter in determining the 21-cm signal. Our main results are shown in Figs. 1-2.
In Fig. 2, the $z\lesssim 10$ error bars indicate the expected noise level from HERA in its design sensitivity under the conservative assumption of a moderate foreground scenario, evaluated with 21cmSense Pober et al. (2013, 2014), while the $z=14$ error bar illustrates the sensitivity expected for SKA, estimated by Ref. Meriot and Semelin (2024). This figure demonstrates that the effects discussed in this work should matter in future analyses of 21-cm signal measurements.
This paper is organized as follows. In Sec. II, we examine the impact of using the baryon density in the computation of the 21-cm signal during the dark ages. We then move to analyzing the role of the baryons and CDM during cosmic dawn and reionization in Sec. III. We discuss in Sec. IV our results from the previous section and write our conclusions in Sec. V. In addition, this paper contains several appendices. In Appendix A, we study the sensitivity of the 21-cm signal to the precise evolution of CDM and the total matter density fields. In Appendix B, we study the Boltzmann equation for the spin temperature. In Appendix C, we verify the output of 21cmFirstCLASS, by comparing the temperature fluctuations at the dark ages with the analytical prediction from linear perturbation theory. Finally, we derive the equations for second-order scale-dependent LPT in Appendix D.
Throughout this work, we adopt cosmological parameters from Planck18 best-fit values Aghanim et al. (2020), without including data from baryon acoustic oscillations (BAO), namely a Hubble constant of $h=0.6736$, a total matter and baryon density parameters of $\Omega_{m}=0.3153$ and $\Omega_{b}=0.0493$, a primordial curvature power spectrum with amplitude $A_{s}=2.1\times 10^{-9}$ and spectral index $n_{s}=0.9649$, and an optical depth to reionization [^2] While 21cmFirstCLASS is capable of computing $\tau_{\mathrm{reio}}$ from the simulation in a self-consistent manner (see Refs. Shmueli et al. (2023); Montero-Camacho et al. (2024)), the exact value of $\tau_{\mathrm{reio}}$ is not crucial for this work as it affects only the CMB power spectrum. of $\tau_{\mathrm{reio}}=0.0544$. We also conservatively assume two massless neutrinos with a third massive neutrino of mass $0.06\,\mathrm{eV}$. For the astrophysical parameters, we work with the fiducial values of EOS2021 (see Table 1 in Ref. Muñoz et al. (2022)), but we assume $L_{X}/\mathrm{SFR}=10^{40}\,\mathrm{erg\cdot yr\cdot sec^{-1}}\cdot M_{\odot}^{-1}$ for both popII and popIII stars. All of our 21cmFirstCLASS simulations are realized with a box of size $400\,\mathrm{Mpc}$ and a simulation cell of $2\,\mathrm{Mpc}$. All the formulae in this paper are expressed in the CGS unit system. To reduce clutter, we often do not write explicitly the independent arguments of the quantities in our equations (e.g. position, redshift, wavenumber, etc) and they should be inferred from the context.
II Dark ages
II.1 Matter density fields at the dark ages
According to the standard cosmological model, two kinds of matter controlled the evolution of the Universe at early times, before the first atoms were formed. The first kind are baryons, which include hydrogen and helium nuclei and electrons and protons that interact strongly with photons due to their electric charge. The second kind is collisionless CDM. While the baryons were subject to radiation pressure on small scales, CDM was able to freely cluster under the influence of self-gravitational forces. As the Universe expanded, it eventually reached a low enough temperature such that atoms could have survived immediate ionization for the first time. This promptly led to a dramatic increase in the photon mean free path, thus releasing the freely propagating cosmic microwave background (CMB) radiation. This key moment in the history of the Universe, dated at redshift $z\approx 1100$, is known as “recombination”.
Figure 3: Left panel: the scale-dependent growth factor for baryons; note the fluctuations at very high $z$, where the baryon fluid oscillates being still partially coupled to photons. Right panel: comparison with the scale-independent growth factor.
During the dark ages, the fluctuations in the matter density field were still much smaller than unity and linear perturbation theory can be applied in that epoch. After recombination took place, on sub-horizon scales larger than the Jeans scale, the linearized continuity and Euler equations for the baryon and CDM fluids are identical Baumann (2022); Dodelson (2003),
$\dot{\delta}_{b}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{b}$ $=$ 00 0 (1)
$\dot{\mathbf{v}}_{b}+H\mathbf{v}_{b}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ 00 0 (2)
$\dot{\delta}_{c}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{c}$ $=$ 00 0 (3)
$\dot{\mathbf{v}}_{c}+H\mathbf{v}_{c}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ $0,$ (4)
where $\delta_{b}$ ($\delta_{c}$) is the baryon (CDM) energy over-density, $\mathbf{v}_{b}$ ($\mathbf{v}_{c}$) is the proper baryon (CDM) peculiar velocity, $a$ is the scale factor, and $H\equiv\dot{a}/a$ is the Hubble parameter. Overdots denote derivatives with respect to the cosmological time $t$. Eqs. (1)-(4) are accompanied with the Poisson equation for the gravitational potential $\Phi$,
$\nabla^{2}\Phi=4\pi G\bar{\rho}_{m}a^{2}\delta_{m},$ (5)
where $G$ is Newton’s gravitational constant, $\bar{\rho}_{m}$ is the mean matter energy density and $\delta_{m}$ the matter energy over-density,
$\delta_{m}=\Omega_{m}^{-1}\left(\Omega_{c}\delta_{c}+\Omega_{b}\delta_{b}\right).$ (6)
We can combine Eqs. (1)-(6) into two coupled second-order differential equations of the same form ($i=b,c$ for baryons and CDM),
$\ddot{\delta}_{i}+2H\dot{\delta}_{i}-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}=0.$ (7)
Even though the equations of motion for $\delta_{b}$ and $\delta_{c}$ are the same (at the relevant scales), their solutions are different because of the initial conditions at the time of recombination; prior that moment, CDM has already clustered on all scales, while the energy density of the baryons was oscillating due to the strong coupling with CMB photons. Thus, $\delta_{b}\leq\delta_{c}$ on all scales.
If $\delta_{b}=\delta_{c}=\delta_{m}$, then Eq. (7) becomes a single second order differential equation for $\delta_{m}$. Since this differential equation is the same on all scales, variable separation can be performed, $\delta_{m}\left(\mathbf{x},z\right)=\delta_{0}\left(\mathbf{x}\right)D\left(z\right)$, where $\delta_{0}\left(\mathbf{x}\right)$ is a function of only the comoving coordinate $\mathbf{x}$ in real space, and $D\left(z\right)$ is the scale-independent growth factor (SIGF), a function of redshift only that obeys the same differential equation for $\delta_{m}$,
$\ddot{D}+2H\dot{D}-4\pi G\bar{\rho}_{m}D=0,$ (8)
with the boundary conditions $D\left(z=0\right)=1$ and $D\left(z\to\infty\right)=0$. Because $\delta_{c}\geq\delta_{b}$ and $\Omega_{c}>\Omega_{b}$, it is often approximated $\delta_{m}\approx\delta_{c}$ and $\delta_{c}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)D\left(z\right)$ in Fourier space. In this work we use a more precise expression for the baryon and CDM over-densities, [^3] Here we assume that $\delta_{b}\approx\delta_{c}=\delta_{0}$ at $z=0$, which is very well justified as at this redshift the baryons have already clustered to the gravitational potential wells sourced by CDM.
$\delta_{i}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)\mathcal{D}_{i}\left(k,z\right)\qquad i=b,c.$ (9)
where $\mathcal{D}_{i}\left(k,z\right)$ is the scale-dependent growth factor (SDGF) for species $i$. By construction, it is defined as
$\mathcal{D}_{i}\left(k,z\right)\equiv\frac{\mathcal{T}_{i}\left(k,z\right)}{\mathcal{T}_{i}\left(k,z=0\right)}\qquad i=b,c.$ (10)
where $\mathcal{T}_{i}\left(k,z\right)=\delta_{i}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the matter density transfer function of species $i$ and $\mathcal{R}$ is the primordial curvature perturbation.
The CDM SDGF is nearly scale-invariant and can be very well approximated by $\mathcal{D}_{c}\left(k,z\right)\approx D\left(z\right)$, apart for BAO features that appear at $z\lesssim 1$ (while adopting the SIGF for the CDM density field causes “fake BAOs” at high redshifts, see more details in Appendix A). In contrast, the baryon SDGF $\mathcal{D}_{b}\left(k,z\right)$ is much smaller than $D\left(z\right)$ at high redshifts, while at low redshifts they are nearly the same due to the increasing coupling due to gravitation between the baryon and CDM fluids. This is demonstrated in Fig. 3, where BAO features are clearly manifest. Since $\mathcal{D}_{b}\left(k,z\right)\ll\mathcal{D}_{c}\left(k,z\right)$ at the dark ages, we have $\delta_{b}\left(\mathbf{x},z\right)\ll\delta_{c}\left(\mathbf{x},z\right)$ from Eq. (9), as illustrated in Fig 4. This has some important implications on the 21-cm signal at that epoch.
Figure 4: Left (right) panel: realization of the CDM (baryon) over-density field at $z=300$.
II.2 Brightness temperature at the dark ages
The quantity of most interest in the study of the 21-cm signal is the brightness temperature relative to the CMB, defined as
$T_{21}=\frac{T_{s}-T_{\gamma}}{1+z}\left(1-\mathrm{e}^{-\tau_{21}}\right).$ (11)
Here, $T_{\gamma}$ is the characteristic temperature of the radio background, which we conservatively assume to follow the CMB temperature, $T_{\gamma}\propto\left(1+z\right)$. We consider this background to be homogeneous in space on the scales of interest for 21-cm tomography, $T_{\gamma}=\bar{T}_{\gamma}$ (see however studies that included background sources other than the CMB Feng and Holder (2018); Fialkov and Barkana (2019); Reis et al. (2020); Sikder et al. (2023)). The 21-cm optical depth $\tau_{21}$ is given by
$\tau_{21}=\frac{3\hbar A_{10}c\lambda_{21}^{2}n_{\mathrm{HI}}}{16Hk_{B}T_{s}}\left(1+\frac{1}{aH}\frac{d\left(\hat{\mathbf{n}}\cdot\mathbf{v}_{b}\right)}{d\left(\hat{\mathbf{n}}\cdot\mathbf{x}\right)}\right)^{-1},$ (12)
[…]
The quantity $T_{s}$ that appears in Eqs. (11)-(12) is the spin temperature. This quantity is defined as $n_{1}/n_{0}\equiv 3\,\mathrm{exp}\left(-T_{\star}/T_{s}\right)$, where $n_{1}$ ($n_{0}$) denotes the number density of hydrogen atoms in the hyperfine triplet (singlet) state and $T_{\star}\equiv\hbar c/\left(k_{B}\lambda_{21}\right)\approx 68.2\,\mathrm{mK}$. In theory, the solution for $T_{s}$ has to be obtained by solving the Boltzmann equation for $n_{1}$ and $n_{0}$. Yet, we show in Appendix B that the solution for this equation is very well approximated by the well known expression for $T_{s}$ in equilibrium, even in the dark ages where $H\gg A_{10}$. This equilibrium expression reads
$T_{s}^{-1}=\frac{x_{\mathrm{CMB}}T_{\gamma}^{-1}+x_{\mathrm{coll}}T_{k}^{-1}+\tilde{x}_{\alpha}T_{\alpha}^{-1}}{x_{\mathrm{CMB}}+x_{\mathrm{coll}}+\tilde{x}_{\alpha}},$ (13)
[…]
In order to compute $T_{s}$ one has to solve the following differential equation for $T_{k}$ that arises from the first law of thermodynamics,
$\frac{dT_{k}}{dt}=$ $-2HT_{k}+\frac{2}{3}\frac{T_{k}}{1+\delta_{b}}\frac{d\delta_{b}}{dt}-\frac{T_{k}}{1+x_{e}}\frac{dx_{e}}{dt}$ (14)
$+\left.\frac{dT_{k}}{dt}\right|_{\mathrm{CMB}}+\left.\frac{dT_{k}}{dt}\right|_{X},$
where $x_{e}\equiv n_{e}/\left(n_{\mathrm{H}}+n_{\mathrm{He}}\right)$ is the free-electron fraction, with $n_{e}$, $n_{\mathrm{H}}$ and $n_{\mathrm{He}}$ the free electron, hydrogen nuclei and helium nuclei number densities, respectively. The first line in Eq. (14) represents the change in temperature in the absence of external heating sources; the first term accounts for adiabatic cooling due to the expansion of the Universe, the second term accounts for adiabatic heating due to clustering of matter and the third term accounts for adiabatic cooling/heating as a result of increasing/lowering the total number of baryons in reionization/recombination. The second line in Eq. (14) represents heating by external sources, with the first (second) term accounting for heating by CMB (X-ray) photons. The expression for Compton heating by CMB photons is well-known and can be found in many papers in the literature, including Refs. Peebles (2020); Ma and Bertschinger (1995); Seager et al. (1999, 2000); Barkana and Loeb (2005b); Naoz and Barkana (2005). In contrast, the calculation of X-ray heating, relevant only during the cosmic dawn and reionization epochs, is much more complicated and model dependent, see e.g. Ref. Mesinger et al. (2011).
Another differential equation that has to be solved simultaneously with Eq. (14) is the Boltzmann equation for the free electron fraction $x_{e}$. The simpler form of this equation can be found in Refs. Flitter and Kovetz (2024a, b). 21cmFirstCLASS involves the state of the art code HYREC Ali-Haimoud and Hirata (2011); Lee and Ali-Haïmoud (2020) that solves with great precision the fully-detailed Boltzmann equation for $x_{e}$ while accounting for helium recombination and high-order energy excitations.
During the dark ages, the only relevant matter species that contributes to the 21-cm signal are the baryons, as they source the emission and absorption of 21-cm radiation by hydrogen atoms, the recombination rate that governs $T_{k}$, and the coupling of $T_{s}$ to $T_{k}$ via collisional excitations. Thus the fluctuations in the 21-cm signal during the dark ages are entirely proportional to $\delta_{b}$. Because $\delta_{b}\ll 1$ during the dark ages, the scale-dependent evolution of the baryon density field barely affects the global 21-cm signal, i.e. the sky-averaged brightness temperature as a function of redshift. This is demonstrated in Fig. 5, where the baryon over-density field is evolved with the SIGF (blue curve), or the SDGF (orange curve).
Figure 5: The 21-cm global signal as a function of redshift, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). The curves almost completely overlap each other.
Although the scale-dependent evolution of $\delta_{b}$ does not affect the global 21-cm signal, it does affect its two-point statistics. Specifically, let us consider the 21-cm power spectrum,
$\Delta_{21}^{2}\left(k,z\right)=\frac{k^{3}\bar{T}_{21}^{2}\left(z\right)P_{21}\left(k,z\right)}{2\pi^{2}},$ (15)
where $\bar{T}_{21}$ is the global brightness temperature and $P_{21}\left(k,z\right)$ is the angle-averaged Fourier transform of the two-point correlation function $\langle\delta_{21}\left(\mathbf{x},z\right)\delta_{21}\left(\mathbf{x}^{\prime},z\right)\rangle$, while $\delta_{21}$ is the local contrast in the brightness temperature, $\delta_{21}\left(\mathbf{x},z\right)\equiv T_{21}\left(\mathbf{x},z\right)/\bar{T}_{21}\left(z\right)-1$. In a previous work Flitter and Kovetz (2024b), we have shown that during the dark ages $\Delta_{21}^{2}\left(k,z\right)\propto\mathcal{T}_{b}^{2}\left(k,z\right)\ll\mathcal{T}_{c}^{2}\left(k,z\right)$. The analysis in that work was based on the same SDGF formalism we present here, which we have shown to agree well with fully relativistic Boltzmann-Einstein solvers. In Fig. 6, we show $\Delta_{21}^{2}\left(k,z\right)$, as calculated by 21cmFirstCLASS after the incorporation of the SDGF in the code (we verify the correctness of our results in Appendix C). As expected, $\Delta_{21}^{2}\left(k,z\right)$ is suppressed during the dark ages when $\delta_{b}$ is correctly evolved with the SDGF.
Figure 6: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). For a zoomed-in version of the figure at cosmic dawn, see Fig. 15 in Appendix D.
III Cosmic dawn and reionization
III.1 Baryons
The role of the baryons in shaping the 21-cm signal is enhanced when the dark ages are over. As dark matter halos collapse, the gravitational pull that the baryons experience is increased. If the cooling timescale is shorter than the free-fall timescale, zero-metalicity popIII stars can form via gas accretion, mostly in molecular cooling galaxies (MCGs) Tegmark et al. (1997); Abel et al. (2002); Bromm and Larson (2004); Haiman and Bryan (2006); Trenti (2010); Plombat and Puy (2024). Baryonic feedbacks, such as supernovae Springel and Hernquist (2003); Wise and Abel (2008); Pawlik and Schaye (2009), photo-heating Hui and Gnedin (1997); Okamoto et al. (2008); Ocvirk et al. (2020); Katz et al. (2020) and Lyman-Werner radiation Bromm and Larson (2004); Fialkov et al. (2013); Haiman et al. (1997) hamper the star formation rate (SFR). Still, stars continue to form in more massive halos that host atomic cooling galaxies (ACGs) Qin et al. (2020, 2021a); Krumholz (2012). Those halos typically are assumed to contain mainly popII stars with low metalicity.
[…]
The local SFRD does not depend only on the local baryon density, but also on local velocities, specifically on the relative velocity between baryons and CDM, $\mathbf{v}_{cb}=\mathbf{v}_{c}-\mathbf{v}_{b}$. From Eqs. (2) and (4), we see that $\dot{\mathbf{v}}_{cb}=-H\mathbf{v}_{cb}$, which has the temporal solution $\mathbf{v}_{cb}\left(\mathbf{x},z\right)\propto\left(1+z\right)$. As for the spatial solution, adiabatic initial conditions imply Gaussian distributions for each one of the three components of $\mathbf{v}_{cb}$, and a Maxwell-Boltzmann distribution for $v_{cb}\equiv\left(\mathbf{v}_{cb}\cdot\mathbf{v}_{cb}\right)^{1/2}$. The variance of this distribution can be determined by defining $\theta_{cb}\equiv a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{cb}$, and then
$\langle v_{cb}^{2}\left(z\right)\rangle=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\frac{\mathcal{T}^{2}_{\theta_{cb}}\left(k,z\right)}{k^{2}},$ (16)
[…]
Figure 7: The effect of the baryon-CDM relative velocity $v_{cb}$ on the 21-cm global signal. In the green curve, the effect of $v_{\mathrm{cb}}$ enters only through the modification of the minimum halo mass that can host popIII stars, as implemented in 21cmFirstCLASS. In the orange curve, small-scale suppression effects in $\sigma\left(M_{h},z\right)$ are also included, as in 21cmFAST, see more details in Sec. IV. The small extra delay between the orange and green curves can be seen more clearly in Fig. 1 (as the difference between the solid and dashed green curves at $z\gtrsim$10).
Once galaxies have started to form inside halos, the baryon density field is far from homogeneous as $\delta_{b}>1$ in regions of galaxy clusters. Thus, during cosmic dawn and reionization epochs linear perturbation theory is no longer applicable in determining $\delta_{b}\left(\mathbf{x},z\right)$. Non-linear perturbation theories, such as standard perturbation theory (or effective field theory) Baumann et al. (2012); Carrasco et al. (2012, 2014); Hertzberg (2014); Porto et al. (2014); Senatore (2015); Senatore and Zaldarriaga (2014); Baldauf et al. (2015), renormalized perturbation theory Crocce and Scoccimarro (2006); Bernardeau et al. (2008); Montesano et al. (2010), the renormalization group approach McDonald (2007), closure theory Taruya and Hiramatsu (2008), TimeRG theory Matarrese and Pietroni (2007); Pietroni (2008), or time-sliced perturbation theory Blas et al. (2016), must be used in order to find the “non-linear” perturbation in the baryon density field $\delta_{b}^{\mathrm{NL}}$.
In this work, we consider a 2nd-order Lagrangian perturbation theory Friedman and Schutz (1978); Buchert (1989); Moutarde et al. (1991); Buchert (1992, 1993); Bouchet et al. (1995); Hivon et al. (1995); Buchert (1995); Tatekawa (2005); Matsubara (2008); Carlson et al. (2009); Zheligovsky and Frisch (2014) (2LPT), as in 21cmFAST, though we improved its calculations in 21cmFirstCLASS to account for the scale-dependent growth of the baryons. Readers who wish to learn more about the scale-dependent 2LPT can find more details in Appendix D. The importance of evolving the baryon density field non-linearly is evident in Fig. 8; while linear perturbation theory seems to give a rough estimate for the distribution of baryonic matter on very large scales, it fails to capture the clustering of baryons at low redshifts and small scales, and it wrongly predicts a more homogeneous density field. When computing the 21-cm power spectrum, the imprecise calculation from linear perturbation theory yields an order unity error at $z\lesssim 8$, as demonstrated in Fig. 9. Importantly, this is the regime where upcoming 21-cm interferometers are mostly sensitive (c.f. Fig. 2).
Figure 8: Left (right) panel: realization of the baryon over-density field at $z=6$ when evolved linearly (non-linearly).
Figure 9: The 21-cm power spectrum as a function of redshift at $k=0.3\,\mathrm{Mpc}^{-1}$, when the baryon and the total matter over-density fields are evolved either linearly or non-linearly. The colors of the curves here match the same colors in Figs. 1-2, namely the blue, orange and green curves can be associated with 21cmFAST, Zeus21-like, and 21cmFirstCLASS, respectively. Here though, unlike Figs. 1-2, in all curves we distinguish between baryons and CDM (i.e. in all curves we use the SDGF), and we have the same $v_{cb}$ delaying effect (see differences in Fig. 7).
III.2 Total matter
During the cosmic dawn and reionization epochs, the 21-cm signal does not depend only on the baryon over-density $\delta_{b}$, but also on the CDM over-density field $\delta_{c}$, or more precisely on the total matter over-density $\delta_{m}$, Eq. (6). This is because the SFRD is given by the following expression Madau et al. (1997); Schneider et al. (2021); Muñoz et al. (2022); Muñoz (2023); Cruz et al. (2024b),
[…]
where $\dot{M}_{*}\left(\mathbf{x},M_{h},z\right)$ is the local stellar mass production rate in a halo of mass $M_{h}$ and $dn\left(\mathbf{x},M_{h},z\right)/dM_{h}$ is the conditional halo mass function (HMF); it is conditional in the sense that it depends on the value of the local total matter over-density $\delta_{m}\left(\mathbf{x},z\right)$. The mean of the conditional HMF is the “standard” HMF, defined as Lukic et al. (2007)
$\frac{d\bar{n}_{h}\left(M_{h},z\right)}{dM_{h}}=\frac{\bar{\rho}_{m,0}}{M_{h}}\frac{d\ln\sigma^{-1}\left(M_{h},z\right)}{dM_{h}}f\left(\sigma\right),$ (18)
where $\bar{\rho}_{m,0}$ is the background value of the current (at $z=0$) total matter density field and $\sigma^{2}\left(M_{h},z\right)$ is the variance of the linear total matter density field, smoothed with a top-hat filter in real space of comoving radius $R_{M_{h}}=\left(3M_{h}/4\pi\bar{\rho}_{m,0}\right)^{1/3}$,
$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z\right)W^{2}\left(kR_{M_{h}}\right),$ (19)
where $\mathcal{T}_{m}\left(k,z\right)\equiv\delta_{m}^{\mathrm{lin}}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the linear matter density transfer function and $W\left(kR\right)=3\left(\sin kR-kR\cos kR\right)/\left(kR\right)^{3}$. The function $f\left(\sigma\right)$ is the HMF fitting function and is calibrated on data from N-body simulations. Perhaps the simplest choice is to rely on the purely Gaussian fitting function from the Press-Schechter (PS) Press and Schechter (1974) HMF. In this work, we use instead the Sheth-Tormen (ST) HMF Sheth and Tormen (1999); Sheth et al. (2001); Sheth and Tormen (2002) (but see other popular mass functions in Refs. Jenkins et al. (2001); Warren et al. (2006); Reed et al. (2007); Tinker et al. (2008); Watson et al. (2013)).
In evaluating the integral in Eq. (17), the following quantity is often considered,
[…]
This is the local collapsed fraction of halos more massive than $M_{\mathrm{min}}\left(z\right)$. The collapsed fraction is a very useful quantity since the extended Press-Schechter (EPS) theory predicts its value analytically Bond et al. (1991). According to EPS, for a PS HMF, in a large-scale region of comoving radius $R$, centered at $\mathbf{x}$, with a smoothed over-density $\delta_{m,R}$, the collapsed fraction of halos above $M_{\mathrm{min}}$ at redshift $z$ is
$f_{\mathrm{coll}}^{\mathrm{EPS}}\left(\mathbf{x},z;M_{\mathrm{min}}\right)=$
$\hskip 71.13188pt\mathrm{erfc}\left(\frac{\delta_{\mathrm{crit}}-\delta_{m,R}\left(\mathbf{x},z\right)}{\sqrt{2\left[\sigma^{2}\left(M_{\mathrm{min}},z\right)-\sigma^{2}\left(M_{R},z\right)\right]}}\right),$ (21)
where $\mathrm{erfc}\left(\cdot\right)$ is the complementary error function, $\delta_{\mathrm{crit}}=1.686$, and
$M_{R}$ $=$ $\frac{4\pi}{3}R^{3}\bar{\rho}_{m,0}$ (22)
$=$ $1.66\times 10^{11}\left(\frac{h^{2}\Omega_{m}}{0.143}\right)\left(\frac{R}{1\,\mathrm{Mpc}}\right)^{3}\,M_{\odot}.$
Note that in writing Eq. (III.2) we have used the redshift-dependent $\sigma\left(M_{h},z\right)$ as given by Eq. (19), without imposing the SIGF approximation, $\sigma\left(M_{h},z\right)\propto D\left(z\right)$, which is less accurate at high redshifts and low masses (see more details in Appendix A).
IV Discussion
One delicate nuance regarding the formalism presented above concerns the assumed matter density field that is used in Eq. (III.2). Press-Schechter theory Sheth and Tormen (1999), and the extended Press-Schechter theory Bond et al. (1991), are both established on linear cosmological perturbation theory. Hence, from a purely theoretical point of view, it would be more consistent to evaluate the smoothed $\delta_{m}$ that appears in Eq. (III.2), as well as $\sigma\left(M_{h},z\right)$, from linear theory. This is for example the underlying assumption in Zeus21, where the “linear” $\delta_{m}^{\mathrm{lin}}$ still results in non-linear fluctuations in the SFRD due to the erfc function in Eq. (III.2), an effect that can be captured by an exponential fit Muñoz (2023). In contrast, 21cmFAST uses the non-linear $\delta_{m}^{\mathrm{NL}}$ (from 2LPT calculations) in Eq. (III.2). The justification for this choice comes from better agreement with radiative transfer simulations at the epoch of reionization Zahn et al. (2011). [^4] We comment that the non-linear density field contains non-Gaussian statistics and therefore non-vanishing bispectrum, leading to a much more complicated $f_{\mathrm{coll}}$, see Eq. (5) in Ref. Lidz et al. (2013). Evidently, because the fluctuations in the brightness temperature are strongly connected to the fluctuations in the SFRD, the choice of working with either $\delta_{m}^{\mathrm{lin}}$ or $\delta_{m}^{\mathrm{NL}}$ has large implications on the resulting 21-cm power spectrum, as can be seen in Fig. 9. Our readers are hence forewarned.
The second nuance that we wish to highlight is the effect that $v_{cb}$ has on the local SFRD. As previously discussed, the local SFRD is suppressed in regions with large relative velocity. In 21cmFAST and 21cmFirstCLASS this effect is modelled in the code by increasing $M_{\mathrm{min}}$ in MCGs Muñoz et al. (2022),
$M_{\mathrm{min}}^{\mathrm{MCG}}\propto\left[1+A_{v_{cb}}\frac{v_{cb}\left(\mathbf{x},z_{\mathrm{rec}}\right)}{\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle^{1/2}}\right]^{\beta_{v_{cb}}},$ (23)
where $A_{v_{cb}}$ and $\beta_{v_{cb}}$ are free parameters, and $\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle$ is the variance of the $v_{cb}$ field at recombination (see Eq. (16)). This modification in $M_{\mathrm{min}}$ results in having VAOs in the 21-cm power spectrum at large scales Muñoz (2019a, b); Sarkar and Kovetz (2023); Zhang et al. (2024). In addition to the above modification, 21cmFAST also modifies the current matter power spectrum by introducing a multiplicative correction factor of the form $1-A_{p}\exp\left[-\left(\ln\left[k/k_{p}\right]\right)^{2}/\left(2\sigma_{p}^{2}\right)\right]$ Muñoz et al. (2022). This factor arises from non-linear $v_{\mathrm{cb}}$ corrections, and was calibrated at $z=20$ with $A_{p}=0.24$, $k_{p}=300\,\mathrm{Mpc}^{-1}$, and $\sigma_{p}=0.9$. 21cmFirstCLASS has the option to remove this factor since it introduces non-linear corrections to $\delta_{m}^{\mathrm{lin}}$ and $\sigma\left(M_{h},z\right)$, while the EPS formalism requires these quantities to be evaluated from linear theory. Because this factor suppresses power at small scales, removing it advances cosmic dawn to slightly higher redshifts, as can be seen in Fig. 1 and 7. A second forewarning due here.
Finally, a third nuance in the calculation arises due to the appearance of the smoothing radius $R$ in the RHS of Eq. (III.2). According to EPS, this $R$ should correspond to the region in which the local collapsed fraction is evaluated. Hence, we argue that the appropriate $R$ to consider should be associated with the simulation’s cell, $R=\left(3/4\pi\right)^{1/3}L_{\mathrm{cell}}$, where $L_{\mathrm{cell}}$ is the cell size of the low-resolution grid (for better precision, the realization of the perturbed box should be generated on a high-resolution grid). Note that the cell size cannot be arbitrarily small as we ought to have $M_{\mathrm{min}}<M_{R}$ for a non-zero local collapsed fraction (see square root in Eq. (III.2)). However, this condition is satisfied for a typical simulation cell of $R\sim L_{\mathrm{cell}}\sim 1\,\mathrm{Mpc}$; according to Eq. (22) such a smoothing radius corresponds to $M_{R}\sim 10^{11}\,M_{\odot}$ and even in ACGs (whose associated $M_{\mathrm{min}}$ is much larger than in MCGs) Barkana and Loeb (2001)
$M_{\mathrm{min}}^{\mathrm{ACG}}\left(z\right)\approx$
$\hskip 14.22636pt2.6\times 10^{8}\,M_{\odot}\left(\frac{T_{\mathrm{vir}}}{1.98\times 10^{4}\,\mathrm{K}}\frac{10}{1+z}\right)^{3/2}\left(\frac{h^{2}\Omega_{m}}{0.143}\right)^{-1/2},$ (24)
where $T_{\mathrm{vir}}\sim 10^{4}\,\mathrm{K}$ is the typical temperature of a virialized halo of mass $M_{\mathrm{min}}$. Therefore, a simulation’s cell of size $\sim 1\,\mathrm{Mpc}$ contains enough mass to form stars in ACGs and MCGs.
[…]
V Conclusions
Building a robust theoretical framework for the 21-cm signal is challenging, especially during the cosmic dawn and reionization epochs, due to the involved complicated astrophysics of galaxy evolution and star formation. While former studies of the 21-cm signal have focused on refining the astrophysical models, they largely glossed over a trivial but crucial aspect of the signal: it is sourced from neutral hydrogen atoms in the IGM between recombination and reionization, and these hydrogen atoms are baryons. Therefore, even though baryons are coupled gravitationally to CDM at low redshifts, the relevant density field to consider is not the CDM density field, but rather the baryon density field. Approximating the baryon density field with the CDM density field when tracing the evolution of the neutral hydrogen atoms in the IGM, leads to misestimation of the 21-cm signal.
In this work, we distinguished between baryons and CDM to carefully study the cosmological impact of the baryons on the 21-cm signal. Unlike most studies which have adopted the SIGF approximation, whereby $\delta_{b}\propto D\left(z\right)$ and evolves in a scale-independent manner, we computed the baryon density field more precisely in 21cmFirstCLASS using the SDGF $\mathcal{D}_{b}\left(k,z\right)$. We have shown in previous work that this method agrees very well with Boltzmann-Einstein solvers at the dark ages when all perturbations can be evolved linearly. Here, we extended this method to the regime when the Universe entered its non-linear phase, including cosmic dawn when galaxies and stars begun to form, as well as reionization.
We started our analysis from the dark ages, and demonstrated that the SIGF approximation fails poorly in that regime. We then moved to the non-linear epochs of cosmic dawn and reionization. While the exact physics in these epochs is still under uncertainty, especially as relates to the efficiency and production rate of stars and galaxies, two facts are certain regarding the cosmological role of the baryons in shaping the 21-cm signal: (1) their relative velocity with respect to CDM delays the signal, and (2) their non-linear clustering cannot be neglected at low redshifts and on small scales. In addition to understanding how the 21-cm signal depends on the baryon over-density $\delta_{b}$, we also discussed the role of the total matter over-density $\delta_{m}$ in the calculation. This enters through the collapsed fraction, Eq. (III.2), as part of the EPS formalism. While it is clear that the 21-cm signal at cosmic dawn and reionization should depend on the “non-linear” $\delta_{b}$ (i.e., $\delta_{b}$ as evaluated by non-linear theory), two common approaches are considered in the literature for the total matter density field $\delta_{m}$. One option is to use the “linear” $\delta_{m}^{\mathrm{lin}}$ since the EPS formalism is based on linear perturbation theory. Alternatively, one can choose to work with the non-linear $\delta_{m}^{\mathrm{NL}}$, as it was argued that this choice yields a better agreement with radiative transfer simulations. While Zeus21 (21cmFAST) adopts the former (latter) approach, 21cmFirtCLASS enables the user to decide which $\delta_{m}$ should be used in Eq. (III.2). Regardless, as was discussed in the previous section, all codes still make in their calculations some assumptions whose validity should be further examined in the future.
Efforts to measure the cosmological 21-cm signal are being carried out by various experiments. The expected data from these experiments, together with new photometric images from the James Webb Space Telescope Gardner et al. (2006); Mirocha and Furlanetto (2019); Harikane et al. (2023); Naidu et al. (2022); Castellano et al. (2022); Adams et al. (2023); Yan et al. (2023); Hassan et al. (2023), present an exciting opportunity to probe uncharted regions of our Universe, thereby deepening our understanding of the physical processes that governed its early evolution. Such an ambitious goal cannot be fulfilled without having robust theoretical predictions to be matched with the data. We thus believe that this work may serve to bring us one step closer towards this goal.
Acknowledgements.
It is our pleasure to thank Julian B. Muñoz for reviewing the manuscript and sharing his comments. We would also like to thank Andrei Mesinger and Tal Adi for useful discussions. We also acknowledge the efforts of the 21cmFAST and CLASS authors to produce state-of-the-art public 21-cm and CMB codes. JF is supported by the Zin fellowship awarded by the BGU Kreitmann School. SL is supported by an Azrieli International Postdoctoral Fellowship. EDK acknowledges joint support from the U.S.-Israel Bi-national Science Foundation (BSF, grant No. 2022743) and the U.S. National Science Foundation (NSF, grant No. 2307354), as well as support from the ISF-NSFC joint research program (grant No. 3156/23).
Appendix A Scale-dependent matter evolution
Even though CDM is considered as a collisionless fluid in $\Lambda$CDM, the evolution of its density fluctuations is not perfectly scale invariant, due to the gravitational coupling with baryons. Indeed, at low redshifts ($z\lesssim 1$) $\delta_{c}$ and $\delta_{b}$ approach each other, and BAO features from $\delta_{b}$ can be seen in the CDM transfer function $\mathcal{T}_{c}\left(k,z\right)$. This is no longer true at the high redshifts relevant for the 21-cm signal. Thus, the commonly used linear approximation of $\delta_{c}\left(\mathbf{k},z\right)\approx D\left(z\right)\delta_{0}\left(\mathbf{k}\right)$ results in fake BAO features in $\delta_{c}\left(\mathbf{k},z\right)$ at high redshifts ($z\gtrsim 1$). This can be remedied by using the appropriate SDGF for CDM $\delta_{c}\left(\mathbf{k},z\right)\approx\mathcal{D}_{c}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where $\mathcal{D}_{c}\left(k,z\right)$ is defined in Eq. (10). We compare between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ in Fig. 10. From the left panel, it indeed appears as if $\mathcal{D}_{c}\left(k,z\right)$ is scale-invariant, unlike $\mathcal{D}_{b}\left(k,z\right)$ (c.f. Fig. 3). However, a closer comparison between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ on the right panel shows that the two can differ by more than 2% at high redshifts ($z\gtrsim 25$), especially at small scales ($k\gtrsim 1\,\mathrm{Mpc}^{-1}$). Moreover, the fake BAO that the SIGF approximation yields is clearly seen on the right panel, through the non-monotonous behavior of $\mathcal{D}_{c}\left(k,z\right)/D\left(z\right)$ when $k$ is increased.
Figure 10: Left panel: the scale-dependent growth factor for CDM. Right panel: comparison with the scale-independent growth factor. The oscillations with scale on the right panel indicate fake BAOs feature when the SIGF approximation is used.
Figure 11: Top panels: The ratio of $\sigma\left(M_{h},z\right)$ and the Sheth-Tormen HMF, when they are calculated with the SIGF and the SDGF. Bottom left panel: The ratio of the collapsed fraction of halos that host ACGs (with $M_{\mathrm{min}}$ given by Eq. (IV)), when calculated with the SIGF and the SDGF. Bottom right panel: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the matter over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
As we discussed in Sec. III.2, the local SFRD can be computed with the EPS formalism, where the local linear matter over-density $\delta_{m}^{\mathrm{lin}}\left(\mathbf{x},z\right)$ is considered. This quantity can be computed in Fourier space via $\delta^{\mathrm{lin}}_{m}\left(\mathbf{k},z\right)\approx\mathcal{D}_{m}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where
$\mathcal{D}_{m}\left(k,z\right)\equiv\frac{\mathcal{T}_{m}\left(k,z\right)}{\mathcal{T}_{m}\left(k,z=0\right)}\approx\frac{\Omega_{b}}{\Omega_{m}}\frac{\mathcal{T}_{b}\left(k,z\right)}{\mathcal{T}_{b}\left(k,z=0\right)}+\frac{\Omega_{c}}{\Omega_{m}}\frac{\mathcal{T}_{c}\left(k,z\right)}{\mathcal{T}_{c}\left(k,z=0\right)}=\frac{\Omega_{b}}{\Omega_{m}}\mathcal{D}_{b}\left(k,z\right)+\frac{\Omega_{c}}{\Omega_{m}}\mathcal{D}_{c}\left(k,z\right),$ (25)
where we used the excellent approximation that $\mathcal{T}_{b}$, $\mathcal{T}_{c}$ and $\mathcal{T}_{m}$ are all the same at $z=0$. This definition allows us to write $\sigma^{2}\left(M_{h},z\right)$ from Eq. (19) as
$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z=0\right)\mathcal{D}_{m}^{2}\left(k,z\right)W_{\mathrm{ball}}^{2}\left(kR_{M_{h}}\right).$ (26)
In the SIGF approximation, it is assumed that $\mathcal{D}_{m}\left(k,z\right)\approx D\left(z\right)$ and hence $\sigma\left(M_{h},z\right)\propto D\left(z\right)$. Therefore, the fake BAOs induced by the SIGF approximation at high redshifts cause an error of a few percent in $\sigma\left(M_{h},z\right)$, mainly in low halo masses, and an order unity error in the derived HMF, as we show in the top panels of Fig. 11. In making the comparison for the HMF though, it is important to bare in mind that the ST parameters were fitted to the HMF that is computed with the SIGF approximation. It is thus intriguing to wonder if the output from N-body simulations can be better fitted by relaxing the SIGF approximation.
When the collapsed fraction in ACGs is considered, the lower limit of the integral in Eq. (20) is at $M_{h}\sim 10^{8}\,M_{\odot}$, which is coincidently the regime where the error in the HMF is minimum, while higher halo masses less contribute to the integral. Thus, the error in the collapsed fraction of halos that host ACGs is sub-percent, as shown in the bottom left panel in Fig. 11. When MCGs are considered too, their $M_{\mathrm{min}}$ is smaller than the equivalent quantity in ACGs by an order of magnitude, and it is subject to inhomogeneous baryonic feedback such as LW radiation and relative velocity with respect to CDM. We compare in the last panel of Fig. 11 how the SIGF approximation, applied both in ACGs and MCGs, affects the 21-cm power spectrum. As it appears, the error introduced by the SIGF approximation in the HMF at high redshifts has almost no effect on the 21-cm power spectrum. This conclusion however relies on the assumed astrophysical model and the fiducial astrophysical parameters we have chosen. In other models, where the star production rate is predicted to be more efficient at high redshifts, the SIGF approximation may yield larger errors in the 21-cm signal. We defer this investigation for future work.
Appendix B Boltzmann equation for the spin temperature
Eq. (13) is often used in the literature with the justification that the hyperfine transition rate is greater than the expansion rate of the Universe. This is indeed true during cosmic dawn and reionization epochs since
$\frac{H\left(z\right)}{A_{10}}=4.3\times 10^{-4}\left(\frac{h^{2}\Omega_{m}}{0.143}\right)^{1/2}\left(1+z\right)^{3/2}.$ (27)
However, above $z\gtrsim 175$ this fraction becomes larger than unity, so it is unclear whether the equilibrium expression of Eq. (13) is still valid at the dark ages. Indeed, Ref. Lewis and Challinor (2007) solved the fully relativistic Boltzmann Equation for the background spin temperature, while taking optical depth corrections to first order. In this subsection we quantify the validity of Eq. (13), while taking all orders in $\tau_{21}$.
We begin with the definition of the spin temperature from the fraction of hydrogen atoms that occupy the hyperfine triplet state, $n_{1}/n_{0}\equiv\left(g_{1}/g_{0}\right)\exp\left(-T_{*}/T_{s}\right)\approx 3-3T_{*}/T_{s}$, where $g_{1}=3$ ($g_{0}=1$) is the total number of different triplet (singlet) states, and we assume $T_{s}\gg T_{*}$, a very good approximation when Eq. (13) is considered (verified below).
The homogeneous Boltzmann equation for $n_{0}$ (i.e. with no collisional terms) can be determined from the number conservation of hydrogen nuclei $a^{3}n_{\mathrm{H}}=a^{3}n_{\mathrm{HI}}/\left(1-x_{e}\right)$, where $x_{e}\equiv n_{e}/n_{\mathrm{H}}=1-n_{\mathrm{HI}}/n_{\mathrm{H}}$ is the free electron fraction.00footnotetext: For better comparison with Ref. Lewis and Challinor (2007), in this appendix we use a slightly different definition for $x_{e}$ than the main text.
$\frac{d}{dt}\left(a^{3}n_{\mathrm{H}}\right)=\frac{4a^{3}}{1-x_{e}}\left(\frac{dn_{0}}{dt}+3Hn_{0}+\frac{n_{0}}{1-x_{e}}\frac{dx_{e}}{dt}\right)=0,$ (28)
where we approximated $n_{\mathrm{HI}}\approx 4n_{0}$. Without collisional terms, the right hand side is zero, while once we add them,
$\frac{dn_{0}}{dt}+3Hn_{0}+\frac{n_{0}}{1-x_{e}}\frac{dx_{e}}{dt}=-n_{0}\left(C_{01}+P_{01}+B_{01}I_{\mathrm{CMB}}\zeta_{\mathrm{CMB}}\right)+n_{1}\left(C_{10}+P_{10}+A_{10}+B_{10}I_{\mathrm{CMB}}\zeta_{\mathrm{CMB}}\right),$ (29)
[…]
$f_{21}\left(\nu\right)=T_{*}^{-1}\left[T_{s}+\left(T_{\gamma}-T_{s}\right)\mathrm{e}^{-\tau_{21}\chi_{21}\left(\nu\right)}\right],$ (30)
where $\chi_{21}\left(\nu\right)$ is the cumulative function of the 21-cm line profile ($\chi_{21}$ goes from zero on the blue side to unity on the red side). Then, we find that
[…]
where Venumadhav et al. (2018)
$x_{\mathrm{CMB}}\equiv\frac{1-\mathrm{e}^{-\tau_{21}}}{\tau_{21}}.$ (32)
In most places in the literature, e.g., Ref. Loeb and Zaldarriaga (2004), the zeroth order of the optical depth is considered with $x_{\mathrm{CMB}}=1$, and $\zeta_{\mathrm{CMB}}=1$. In Ref. Lewis and Challinor (2007) the first order is considered with $x_{\mathrm{CMB}}=1-\tau_{21}/2$ and $\zeta_{\mathrm{CMB}}=1-\left(1-T_{s}/T_{\gamma}\right)\tau_{21}/2$. Here, instead, we take all orders of the optical depth with $x_{\mathrm{CMB}}$ given by Eq. (32).
In order to proceed with the derivation, we shall assume detailed balance, implying that all sub-process are in equilibrium with their corresponding reverse sub-process. This implies the following relations,
$B_{10}=\frac{\lambda_{21}^{3}}{2\hbar c}A_{10},$ $\quad B_{01}$ $=\frac{g_{1}}{g_{0}}B_{10}=3B_{10},\quad$ (33)
$C_{01}=\frac{g_{1}}{g_{0}}C_{10}\mathrm{e}^{-T_{*}/T_{k}}\approx 3\left(1-\frac{T_{*}}{T_{k}}\right)C_{10},$ $P_{01}=\frac{g_{1}}{g_{0}}P_{10}\mathrm{e}^{-T_{*}/T_{k}}\approx 3\left(1-\frac{T_{*}}{T_{\alpha}}\right)P_{10}.$ (34)
Inserting those relations in Eq. (29) then yields
$\frac{1}{n_{0}}\frac{dn_{0}}{dt}+3H+\frac{1}{1-x_{e}}\frac{dx_{e}}{dt}=\left(\frac{n_{1}}{n_{0}}-3+3\frac{T_{*}}{T_{k}}\right)C_{10}+\left(\frac{n_{1}}{n_{0}}-3+3\frac{T_{*}}{T_{\alpha}}\right)P_{10}$
$+\left[\frac{n_{1}}{n_{0}}\left(1+\frac{T_{\gamma}}{T_{*}}\zeta_{\mathrm{CMB}}\right)-3\frac{T_{\gamma}}{T_{*}}\zeta_{\mathrm{CMB}}\right]A_{10}.$ (35)
We would like to convert this equation to a differential equation for $T_{s}$. For this purpose, we compute the time derivative of $n_{1}/n_{0}$. Using again the fact that $d\left(a^{3}n_{\mathrm{H}}\right)/dt=0$, it is straightforward to show that
$\frac{d}{dt}\left(\frac{n_{1}}{n_{0}}\right)=-\left(\frac{1}{n_{0}}\frac{dn_{0}}{dt}+3H+\frac{1}{1-x_{e}}\frac{dx_{e}}{dt}\right)\left(1+\frac{n_{1}}{n_{0}}\right).$ (36)
On the other hand, since $n_{1}/n_{0}\equiv 3\exp\left(-T_{*}/T_{s}\right)\approx 3-3T_{*}/T_{s}$, then
$\frac{d}{dt}\left(\frac{n_{1}}{n_{0}}\right)=-3T_{*}\frac{dT_{s}^{-1}}{dt}.$ (37)
Combining Eqs. (B)-(37) thus yields
$\frac{dT_{s}^{-1}}{dt}=4\left(T_{k}^{-1}-T_{s}^{-1}\right)C_{10}+4\left(T_{\alpha}^{-1}-T_{s}^{-1}\right)P_{10}+\frac{4}{T_{*}}\left(1-\frac{T_{\gamma}}{T_{s}}\zeta_{\mathrm{CMB}}\right)A_{10}.$ (38)
This is the Boltzmann equation for the (inverse) spin temperature. This equation is almost identical to Eq. (28) in Ref. Lewis and Challinor (2007) with their approximation of $\zeta_{\mathrm{CMB}}\approx 1-\left(1-T_{s}/T_{\gamma}\right)\tau_{21}/2$, but they have an extra $dx_{e}/dt$ term on the LHS which was canceled in our calculations.
To bring this equation to a more familiar form, we make use of $1-\left(T_{\gamma}/T_{s}\right)\zeta_{\mathrm{CMB}}=\left(1-T_{\gamma}/T_{s}\right)x_{\mathrm{CMB}}$ and we plug the dimensionless coupling coefficients, $P_{10}\equiv\left(T_{\gamma}A_{10}/T_{*}\right)\tilde{x}_{\alpha}$ and $C_{10}\equiv\left(T_{\gamma}A_{10}/T_{*}\right)x_{\mathrm{coll}}$. Finally, we also convert the time derivative on the LHS with redshift derivative via $dz/dt=-\left(1+z\right)H\left(z\right)$. We then arrive at
$\frac{dT_{s}^{-1}}{dz}=\frac{4T_{\gamma}A_{10}\left(x_{\mathrm{CMB}}+x_{\mathrm{coll}}+\tilde{x}_{\alpha}\right)}{T_{*}\left(1+z\right)H\left(z\right)}\left[T_{s}^{-1}-\frac{x_{\mathrm{CMB}}T_{\gamma}^{-1}+x_{\mathrm{coll}}T_{k}^{-1}+\tilde{x}_{\alpha}T_{\alpha}^{-1}}{x_{\mathrm{CMB}}+x_{\mathrm{coll}}+\tilde{x}_{\alpha}}\right].$ (39)
From here, we see that if
$\epsilon_{s}\left(z\right)\equiv\left[\frac{4T_{\gamma}A_{10}\left(x_{\mathrm{CMB}}+x_{\mathrm{coll}}+\tilde{x}_{\alpha}\right)}{T_{*}\left(1+z\right)H\left(z\right)}\right]^{-1}$ (40)
is much smaller than unity, then the following solution is an attractor,
$T_{s,\mathrm{eq}}^{-1}\equiv\frac{x_{\mathrm{CMB}}T_{\gamma}^{-1}+x_{\mathrm{coll}}T_{k}^{-1}+x_{\alpha}T_{\alpha}^{-1}}{x_{\mathrm{CMB}}+x_{\mathrm{coll}}+\tilde{x}_{\alpha}}.$ (41)
Then, under this condition ($\epsilon_{s}\ll 1$) we can expand the solution of Eq. (39) to first order in $\epsilon_{s}$, $T_{s}^{-1}=T_{s,\mathrm{eq}}^{-1}+\mathcal{O}\left(\epsilon_{s}\right)$. We plot $\epsilon_{s}\left(z\right)$ during the dark ages in Fig. 12. We find that this quantity reaches maximum at $\epsilon_{s,\max}\left(z=69\right)\approx 10^{-3}$. Thus we estimate
$\frac{\left|T_{s}^{-1}-T_{s,\mathrm{eq}}^{-1}\right|}{T_{s}^{-1}}\lesssim\epsilon_{s,\max}=0.1\%.$ (42)
Hence, solving the Boltzmann equation as in Ref. Lewis and Challinor (2007) is unnecessary, and the equilibrium expression of Eq. (13) holds also at the dark ages with excellent precision.
Figure 12: The inverse dimensionless rate $\epsilon_{s}\left(z\right)$ in the Boltzmann equation for the spin temperature, as given by Eq. (40). We assume that $\tilde{x}_{\alpha}=0$ during the dark ages. During cosmic dawn and reionization epochs, $\tilde{x}_{\alpha}\gg 1$, thereby lowering the value of this quantity even further. Because $\epsilon_{s}\ll 1$, we conclude that the equilibrium expression for $T_{s}$, Eq. (13), is valid also during the dark ages with excellent precision.
Appendix C Temperature fluctuations in the dark ages
In Ref. Flitter and Kovetz (2024b) we discussed the effect of kinetic temperature fluctuations on the brightness temperature fluctuations during the dark ages. Using linear perturbation theory, we showed that the two kinds of fluctuations are approximately related in Fourier space via
$\delta_{21,\mathrm{iso}}\left(k,z\right)\approx\frac{2\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\delta_{b}\left(k,z\right)+\left[\frac{\bar{x}_{\mathrm{CMB}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\frac{\partial\ln\bar{\kappa}_{1-0}^{\mathrm{HH}}}{\partial\ln\bar{T}_{k}}-\frac{1}{1-\bar{T}_{k}/T_{\gamma}}\right]\delta_{T}\left(k,z\right),$ (43)
where $\kappa_{1-0}^{\mathrm{HH}}$ is the collision rate (in units of $\mathrm{cm^{3}/sec}$) of hydrogen atoms with themselves. Here, $\delta_{T}\left(k,z\right)$ and $\delta_{21,\mathrm{iso}}\left(k,z\right)$ are the Fourier transform of the fractional fluctuation in $T_{k}\left(\mathbf{x},z\right)$ and $T_{21}\left(\mathbf{x},z\right)$, respectively, and the latter does not contain fluctuations sourced from peculiar velocity (see Eq. (12)). At the dark ages, the fluctuations $\delta_{T}$ are sourced by the perturbations in the baryon density field $\delta_{b}$, and thus it is useful to define the so called “adiabatic index”,
$c_{T}\left(k,z\right)\equiv\frac{\delta_{T}\left(k,z\right)}{\delta_{b}\left(k,z\right)}.$ (44)
The reason for this name is because for small $\delta_{b}\left(k,z\right)$ we can write $1+\delta_{T}\left(k,z\right)=\left[1+\delta_{b}\left(k,z\right)\right]^{c_{T}\left(k,z\right)}$.
This adiabatic index can be calculated analytically by solving simultaneously the linearly perturbed differential equations for $T_{k}$ and $x_{e}$. We show in Fig. 13 the analytical solution for $c_{T}\left(k,z\right)$ as a function of redshift for $k=0.5\,\mathrm{Mpc}^{-1}$. The blue dashed curve displays the scale-independent $c_{T}\left(z\right)$ which is obtained under the very crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, namely $\delta_{b}\left(k,z\right)$ evolves in a scale-independent manner. In the solid blue curve we have relaxed this assumption and used the exact $\mathcal{D}_{b}\left(k,z\right)$, as defined in Eq. (10). For both curves, we assumed a recombination rate from the RECFAST code Seager et al. (1999, 2000). Evidently, the scale-dependent $c_{T}\left(k,z\right)$ curve reaches higher values than the scale-independent curve.
To verify the consistency of the calculations of 21cmFirstCLASS at the dark ages we also show in Fig. 13 the output of the code for $c_{T}\left(k=0.5\,\mathrm{Mpc}^{-1},z\right)$. This output was achieved by computing the $c_{T}\left(\mathbf{k},z_{i}\right)$ coeval box (in Fourier space) at every redshift iteration $z_{i}$. Because of the scale-dependence, this box is not homogeneous, and so we need to select the appropriate cells that correspond to a desired wavenumber $k_{0}$ (in Fig. 13, $k_{0}=0.5\,\mathrm{Mpc}^{-1}$). We do this selection process using the following criterion,
$k_{0}-\Delta k/2<|\mathbf{k}|<k_{0}+\Delta k/2,$ (45)
where $\Delta k=2\pi/L_{\mathrm{box}}$ and $L_{\mathrm{box}}$ is the box size of the simulation. Once we have the appropriate cells from the $c_{T}\left(\mathbf{k},z_{i}\right)$ box, we take their median, depicted by the orange curve in Fig. 13. There is an excellent agreement between the analytical calculation and the numerical calculation from 21cmFirstCLASS, even though in the former we have assumed the recombination rate from RECFAST, while the latter works with HYREC Ali-Haimoud and Hirata (2011); Lee and Ali-Haïmoud (2020).
Figure 13: The scale-dependent adiabatic index for $k=0.5\,\mathrm{Mpc}^{-1}$ as a function of redshift during the dark ages. The blue curves show the analytic calculation, based on the equations found in Ref. Flitter and Kovetz (2024b), while the orange curve shows the numerical output from 21cmFirstCLASS, see text for details. For the dashed blue curve, we have taken the crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, yielding an inconsistent adiabatic index.
Appendix D Scale-dependent Lagrangian perturbation theory
Normally, in Lagrangian perturbation theory Friedman and Schutz (1978); Buchert (1989); Moutarde et al. (1991); Buchert (1992, 1993); Bouchet et al. (1995); Hivon et al. (1995); Buchert (1995); Tatekawa (2005); Matsubara (2008); Carlson et al. (2009); Zheligovsky and Frisch (2014), the SIGF approximation is taken, namely treating the baryon and CDM density fields the same and solving for the total matter density $\delta_{m}$. In what follows, we relax this approximation and treat the baryons and CDM as two distinguished fluids. In the Lagrangian frame (where we go with the flow, i.e. we follow parcels along their trajectory), the equations of motion are
$\dot{\delta}_{i}+a^{-1}\boldsymbol{\nabla}_{\mathbf{x}}\cdot\mathbf{v}_{i}$ $=$ 00 0 (46)
$\dot{\mathbf{v}}_{i}+H\mathbf{v}_{i}+a^{-1}\boldsymbol{\nabla}_{\mathbf{x}}\Phi$ $=$ 00 0 (47)
where $i=b,c$ for baryons and CDM (to reduce clutter, we omit the $i$ index when we write derivatives with respect to space). Note that we have ignored the sound speed term in the Euler equation as we focus on super Jeans scales. In addition, we have the Poisson equation
$\nabla^{2}\Phi_{\mathbf{x}}=4\pi G\bar{\rho}_{m}a^{2}\delta_{m}=4\pi G\bar{\rho}_{m}a^{2}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}.$ (48)
In LPT, for each species $i$, we relate the initial Lagrangian coordinates $\mathbf{q}^{(i)}$ to the Eulerian coordinates $\mathbf{x}^{(i)}$ via
$\mathbf{x}^{(i)}\left(\mathbf{q}^{(i)},z\right)=\mathbf{q}^{(i)}\left(z_{0}\right)+\mathbf{s}^{(i)}\left(\mathbf{q}^{(i)},z\right),$ (49)
where $\mathbf{s}^{(i)}$ is the displacement field vector for species $i$ and $z_{0}\gg z$ is a high redshift in which the initial conditions for the calculation are given. From here, we see that $\mathbf{v}_{i}\equiv a\dot{\mathbf{x}}^{(i)}=a\dot{\mathbf{s}}^{(i)}$. We now take the divergence of the Euler equation (Eq. (47)), and substitute the Poisson equation (Eq. (48)) to write
$\boldsymbol{\nabla}_{\mathbf{x}}\cdot\left(\ddot{\mathbf{s}}^{(i)}+H\dot{\mathbf{s}}^{(i)}\right)=-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}.$ (50)
We shall focus on irrotational trajectories that obey $\boldsymbol{\nabla}_{\mathbf{q}}\times\mathbf{s}^{(i)}=0$, implying that $\mathbf{s}^{(i)}$ can be expressed as the gradient of a scalar field $S_{i}$,
$\mathbf{s}^{(i)}\left(\mathbf{q},z\right)\equiv\boldsymbol{\nabla}_{\mathbf{q}}S_{i}\left(\mathbf{q},z\right)-\boldsymbol{\nabla}_{\mathbf{q}}S_{i}\left(\mathbf{q},z_{0}\right).$ (51)
Note that by construction, $\mathbf{s}^{(i)}\left(\mathbf{q},z_{0}\right)\equiv 0$, i.e we impose the initial conditions $\mathbf{x}^{(i)}\left(z_{0}\right)\equiv\mathbf{q}^{(i)}\left(z_{0}\right)$.
The continuity equation implies mass conservation for each species $j$, $\rho_{j}\left(\mathbf{x}\right)d^{3}x=\rho_{j}\left(\mathbf{q}\right)d^{3}q$ (here $\rho$ denotes the comoving energy density, without the $\left(1+z\right)^{3}$ scaling). We can use the Jacobian $J_{j}\equiv\det\left(\partial x_{n}^{(j)}/\partial q_{m}\right)$ to write
$\rho_{j}\left(\mathbf{q},z\right)J_{j}d^{3}q=\rho_{j}\left(\mathbf{q},z_{0}\right)d^{3}q\qquad\implies\qquad J_{j}=\frac{\rho_{j}\left(\mathbf{q},z_{0}\right)}{\rho_{j}\left(\mathbf{q},z\right)}\approx\left(1+\delta_{j}\right)^{-1}$ (52)
where the last equality assumes that $\rho_{j}\left(\mathbf{q},z\right)=\bar{\rho}_{j}$, i.e. that the initial density field is nearly unperturbed at $z=z_{0}$. This assumption can be relaxed, although the approximation $J_{j}\approx\left(1+\delta_{j}\right)^{-1}$ will allow us to write Eq. (56) without $S_{i}\left(\mathbf{q},z_{0}\right)$ correction terms. At the end, keeping $S_{i}\left(\mathbf{q},z_{0}\right)$ and $\delta_{i}\left(\mathbf{q},z_{0}\right)$ yields the same solutions we derive, which is why we omit them for brevity. Combining Eqs. (50), (51) and (52) then gives
$\boldsymbol{\nabla}_{\mathbf{x}}\cdot\boldsymbol{\nabla}_{\mathbf{q}}\left(\ddot{S}_{i}+2H\dot{S}_{i}\right)=-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\left(J_{j}^{-1}-1\right).$ (53)
This is one equation we have. We also have an equation that relates the gradients (Einstein summation convention is assumed for the dummy index $m$),
$\frac{\partial}{\partial q_{n}}=\frac{\partial}{\partial x_{n}}+\frac{\partial s_{m}^{(i)}}{\partial q_{n}}\frac{\partial}{\partial x_{m}}=\frac{\partial}{\partial x_{n}}+\frac{\partial^{2}S_{i}}{\partial q_{n}\partial q_{m}}\frac{\partial}{\partial x_{m}}.$ (54)
Finally, we have an equation that relates $J_{j}$ to $S_{j}$. Since for any $3\times 3$ matrix $\mathbf{M}$, we have $\det\mathbf{M}=\epsilon_{ijk}\epsilon_{abc}M_{ia}M_{jb}M_{kc}$, it is straightforward to show that
$J_{j}\equiv\det\left(\frac{\partial x_{n}^{(j)}}{\partial q_{m}}\right)=\det\left(\delta_{nm}+\frac{\partial s_{n}^{(j)}}{\partial q_{m}}\right)=1+\boldsymbol{\nabla}_{\mathbf{q}}\cdot\mathbf{s}^{(j)}+\frac{1}{2}\left[\left(\boldsymbol{\nabla}_{\mathbf{q}}\cdot\mathbf{s}^{(j)}\right)^{2}-\frac{\partial s_{n}^{(j)}}{\partial q_{m}}\frac{\partial s_{m}^{(j)}}{\partial q_{n}}\right]+\det\left(\frac{\partial s_{n}^{(j)}}{\partial q_{m}}\right),$ (55)
or,
$J_{j}=1+\nabla_{\mathbf{q}}^{2}S_{j}+\frac{1}{2}\left[\left(\nabla_{\mathbf{q}}^{2}S_{j}\right)^{2}-\left(\frac{\partial^{2}S_{j}}{\partial q_{n}\partial q_{m}}\right)^{2}\right]+\det\left(\frac{\partial^{2}S_{j}}{\partial q_{n}\partial q_{m}}\right).$ (56)
Equations (53), (54), (56) form a set of non-linear equations for $S_{i}$. We use perturbative approach to solve them.
D.1 First order - Zel’dovich approximation
Let us expand $S_{i}\left(\mathbf{q},z\right)$ to its first order, $S_{i}\left(\mathbf{q},z\right)\approx S_{i}^{(1)}\left(\mathbf{q},z\right)$. Then, to first order,
$J_{j}\approx 1+\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}\implies\delta_{j}=J_{j}^{-1}-1=\left(1+\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}\right)^{-1}-1\approx-\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}.$ (57)
From here, we see that in Fourier space,
$S_{j}^{\left(1\right)}\left(\mathbf{k},z\right)=\frac{\delta_{j}\left(\mathbf{k},z\right)}{k^{2}}\approx\mathcal{D}_{j}\left(k,z\right)\frac{\delta_{0}\left(\mathbf{k}\right)}{k^{2}},$ (58)
where we used the linear expression for $\delta_{i}$ (Eq. (9)) as we seek first order solution to $S_{i}$. To see that this solution is consistent, we expand the Euler-Poisson equation, Eq. (53), to first order,
$\nabla_{\mathbf{q}}^{2}\left(\ddot{S}_{i}^{\left(1\right)}\left(\mathbf{q},z\right)+2H\dot{S}_{i}^{\left(1\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(1\right)}\left(\mathbf{q},z\right)\right)=0.$ (59)
Then, by plugging $\delta_{i}\left(\mathbf{q},z\right)=-\nabla_{\mathbf{q}}^{2}S_{i}^{\left(1\right)}\left(\mathbf{q},z\right)$, we find
$\ddot{\delta}_{i}\left(\mathbf{q},z\right)+2H\dot{\delta}_{i}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}\left(\mathbf{q},z\right)=0,$ (60)
which is indeed the equation of motion for $\delta_{i}$ in linear perturbation theory, above the Jeans scale (c.f. Eq. (7)). Furthermore, the homogeneity assumption we made in Eq. (52) implies that $\delta_{i}$ and $S_{i}^{(1)}$ have the same zero initial conditions at $z=z_{0}$. Thus, the solution to $\mathbf{s}_{i}^{\left(1\right)}$ in Fourier space reads
$\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)=\left[\mathcal{D}_{i}\left(k,z\right)-\mathcal{D}_{i}\left(k,z_{0}\right)\right]\frac{i\mathbf{k}}{k^{2}}\delta_{0}\left(\mathbf{k}\right).$ (61)
This equation is consistent with our ansatz in Ref. Flitter and Kovetz (2024b) that was based on the Zel’dovich approximation Zel’dovich (1970); White (2014); Hidding et al. (2014). Note that given the SDGF, the solution for $\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)$ is independent of its history, e.g. we can find $\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z=20\right)$ before we solve for $\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z=30\right)$. In addition, note that $\mathbf{s}_{i}^{\left(1\right)}$ cannot be separated to its spatial and temporal components, as this separation cannot be done for the SDGF $\mathcal{D}_{i}\left(k,z\right)$. However, this becomes possible when replacing the SDGF with the SIGF $\mathcal{D}_{i}\left(k,z\right)\to D\left(z\right)$,
$\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)\approx\left[D\left(z\right)-D\left(z_{0}\right)\right]\frac{i\mathbf{k}}{k^{2}}\delta_{0}\left(\mathbf{k}\right),$ (62)
as commonly done in the literature Efstathiou et al. (1985); Sirko (2005); Mesinger and Furlanetto (2007).
D.2 Second order - 2LPT
Next, we expand $S_{i}$ to its second order, $S_{i}\left(\mathbf{q},z\right)\approx S_{i}^{(1)}\left(\mathbf{q},z\right)+S_{i}^{(2)}\left(\mathbf{q},z\right)$. The second order expansion of the Jacobian is
$J_{j}\approx 1+\nabla_{\mathbf{q}}^{2}\left(S_{j}^{\left(1\right)}+S_{j}^{\left(2\right)}\right)+\frac{1}{2}\left[\left(\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}\right)^{2}-\left(\frac{\partial^{2}S_{j}^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (63)
Therefore, the second order expansion of the RHS of Eq. (53) is
$-4\pi G\bar{\rho}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\left(J_{j}^{-1}-1\right)\approx 4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\left\{\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}+\nabla_{\mathbf{q}}^{2}S_{j}^{\left(2\right)}-\frac{1}{2}\left[\left(\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}\right)^{2}+\left(\frac{\partial^{2}S_{j}^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right]\right\},$ (64)
while the second order expansion of the LHS is
$\nabla_{\mathbf{x}}\cdot\nabla_{\mathbf{q}}\left(\ddot{S}_{i}+2H\dot{S}_{i}\right)\approx\nabla_{\mathbf{q}}^{2}\left(\ddot{S}_{i}^{\left(1\right)}+\ddot{S}_{i}^{\left(2\right)}+2H\dot{S}_{i}^{\left(1\right)}+2H\dot{S}_{i}^{\left(2\right)}\right)-\frac{\partial^{2}S_{i}^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\frac{\partial^{2}}{\partial q_{n}q_{m}}\left(\ddot{S}_{i}^{\left(1\right)}+2H\dot{S}_{i}^{\left(1\right)}\right).$ (65)
Combining Eqs. (64)-(65), together with the differential equation we obtained for $S^{(1)}$, Eq. (59), yields
$\nabla_{\mathbf{q}}^{2}\left(\ddot{S}_{i}^{\left(2\right)}+2H\dot{S}_{i}^{\left(2\right)}-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(2\right)}\right)=$
$\hskip 56.9055pt\frac{\partial^{2}S_{i}^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\frac{\partial^{2}}{\partial q_{n}q_{m}}\left(\ddot{S}_{i}^{\left(1\right)}+2H\dot{S}_{i}^{\left(1\right)}\right)-2\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\left[\left(\nabla_{\mathbf{q}}^{2}S_{j}^{\left(1\right)}\right)^{2}+\left(\frac{\partial^{2}S_{j}^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (66)
If variable separation was possible, then from Eq. (D.2) we could have obtained two different equations, one for the spatial part of $S_{i}^{(2)}$, and one for its temporal part. However, as we established in Eq. (58), variable separation cannot be performed for $S_{i}^{(1)}$, and thus $S_{i}^{(2)}$ cannot be separated to its spatial and temporal components either.
Note that given $S_{i}^{(1)}$, Eq. (D.2) does not give a differential equation for $S_{i}^{(2)}$, but rather for the Laplacian of $S_{i}^{(2)}$. In order to obtain an integro-differential equation for $S_{i}^{(2)}$, we take the Fourier transform of Eq. (D.2). While applying Fourier transform on the LHS gives a multiplication by $-k^{2}$, taking the Fourier transform of the RHS, being proportional to the square of second derivatives of $S_{i}^{(1)}$, would result in a convolution. We use the fact that in Fourier space $\ddot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)+2H\dot{S}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)=4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(1\right)}\left(\mathbf{k},z\right)$ and apply the inverse Fourier transform to arrive at
$\ddot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)+2H\dot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(2\right)}\left(\mathbf{q},z\right)$
$=-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\int\frac{d^{3}k}{\left(2\pi\right)^{3}}\int\frac{d^{3}k^{\prime}}{\left(2\pi\right)^{3}}\left[\left(\mathbf{k}\cdot\mathbf{k}^{\prime}\right)^{2}S_{i}^{\left(1\right)}\left(\mathbf{k}^{\prime},z\right)-\frac{1}{2}\left[k^{2}k^{\prime 2}+\left(\mathbf{k}\cdot\mathbf{k}^{\prime}\right)^{2}\right]S_{j}^{\left(1\right)}\left(\mathbf{k}^{\prime},z\right)\right]\frac{S_{j}^{\left(1\right)}\left(\mathbf{k},z\right)}{\left|\mathbf{k}+\mathbf{k}^{\prime}\right|^{2}}\mathrm{e}^{i\left(\mathbf{k}+\mathbf{k^{\prime}}\right)\cdot\mathbf{q}}.$ (67)
This integro-differential equation actually contains two equations, one for the baryons ($i=b$) and one for CDM ($i=c$), and these two equations have to be simultaneously solved due to the gravitational coupling between the species, while imposing zero initial conditions for $S_{i}^{\left(2\right)}$ and its derivative at $z=z_{0}$.
D.2.1 Scale independent growth
While in principle Eq. (D.2) can be solved numerically, this is computationally expensive, as for each cell $\mathbf{q}$ in the grid, double integration (that involves all the cells in the box in Fourier space) is required, for both species. To ease the computational cost, we will make the approximation that is done vastly in the literature, of approximating $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$, i.e., we replace the SDGF for each species with a universal SIGF. Then, from Eq. (58), we write
$S_{i}^{\left(1\right)}\left(\mathbf{q},z\right)\approx D\left(z\right)\psi^{\left(1\right)}\left(\mathbf{q}\right)\equiv S^{\left(1\right)}\left(\mathbf{q},z\right),$ (68)
where $\psi^{\left(1\right)}\left(\mathbf{q}\right)$ is the inverse Fourier transform of $\delta_{0}\left(\mathbf{k}\right)/k^{2}$ (or equivalently, in real space $\delta_{0}=-\nabla^{2}_{\mathbf{q}}\psi^{(1)}$). Because $\psi^{(1)}$ is derived from $\delta_{0}$, which is common for all species, under the SIGF approximation the solution to $S_{i}^{\left(1\right)}$ is a universal (i.e. independent of the species $i$) solution $S^{\left(1\right)}$. Plugging this universal solution back to Eq. (D.2) yields
$\ddot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)+2H\dot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(2\right)}\left(\mathbf{q},z\right)$
$\hskip 56.9055pt=4\pi G\bar{\rho}_{m}D\left(z\right)\int\frac{d^{3}k}{\left(2\pi\right)^{3}}\int\frac{d^{3}k^{\prime}}{\left(2\pi\right)^{3}}\left[\frac{1}{2}k^{2}k^{\prime 2}-\frac{1}{2}\left(\mathbf{k}\cdot\mathbf{k}^{\prime}\right)^{2}\right]\frac{\psi^{\left(1\right)}\left(\mathbf{k}\right)\psi^{\left(1\right)}\left(\mathbf{k}^{\prime}\right)}{\left|\mathbf{k}+\mathbf{k}^{\prime}\right|^{2}}\mathrm{e}^{i\left(\mathbf{k}+\mathbf{k^{\prime}}\right)\cdot\mathbf{q}}.$ (69)
Because the RHS of Eq. (D.2.1) is universal for all species, and since $S_{i}^{(2)}$ has zero initial conditions at $z=z_{0}$ for every species $i$, we deduce that the SIGF approximation implies that $S_{i}^{(2)}\left(\mathbf{q},z\right)$ also has a universal solution $S^{(2)}\left(\mathbf{q},z\right)$. This property can be further inferred for any higher order in perturbation theory. Hence, the SIGF approximation corresponds to solving the perturbed equations of motions under the approximation $\delta_{b}\approx\delta_{c}\approx\delta_{m}$, namely we directly solve for the total matter density field.
After substituting the universal solution $S_{i}^{(2)}\left(\mathbf{q},z\right)=S^{(2)}\left(\mathbf{q},z\right)$, Eq. (D.2.1) can be written in the following form,
$\nabla^{2}_{\mathbf{q}}\left[\ddot{S}^{\left(2\right)}\left(\mathbf{q},z\right)+2H\dot{S}^{\left(2\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}S^{\left(2\right)}\left(\mathbf{q},z\right)\right]=-4\pi G\bar{\rho}_{m}D^{2}\left(z\right)\left[\frac{1}{2}\left(\nabla_{\mathbf{q}}^{2}\psi^{\left(1\right)}\right)^{2}-\frac{1}{2}\left(\frac{\partial^{2}\psi^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (70)
Because of the variable separation on the RHS, we see that $S^{(2)}$ can be decomposed to its spatial and temporal parts
$S^{\left(2\right)}\left(\mathbf{q},z\right)=E\left(z\right)\psi^{\left(2\right)}\left(\mathbf{q}\right),$ (71)
where $E\left(z\right)$ and $\psi^{\left(2\right)}\left(\mathbf{q}\right)$ obey the following differential equations,
$\ddot{E}+2H\dot{E}-4\pi G\bar{\rho}_{m}E=-4\pi G\bar{\rho}_{m}D^{2}$ (72)
$\nabla_{\mathbf{q}}^{2}\psi^{\left(2\right)}=\frac{1}{2}\left[\left(\nabla_{\mathbf{q}}^{2}\psi^{\left(1\right)}\right)^{2}-\left(\frac{\partial^{2}\psi^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (73)
As we show below, the solution to $E$ is $E\approx-{3}D^{2}/{7}$ Bouchet et al. (1995); Scoccimarro (1998), and thus the second order correction to the displacement vector is
$\mathbf{s}^{\left(2\right)}\left(\mathbf{q},z\right)=-\frac{3}{7}\left[D^{2}\left(z\right)-D^{2}\left(z_{0}\right)\right]\boldsymbol{\nabla}_{\mathbf{q}}\psi^{\left(2\right)}\left(\mathbf{q}\right).$ (74)
This closed solution for the displacement vector allows one to compute $\mathbf{s}^{\left(2\right)}$ for any $z$, independently from its history.
D.2.2 Solution to $E\left(z\right)$
Before solving for $E\left(z\right)$, we shall solve first the differential equation for the SIGF $D\left(z\right)$. This equation is obtained from Eq. (60) when substituting $\delta_{i}\left(\mathbf{q},z\right)=D\left(z\right)\delta_{0}\left(\mathbf{q}\right)$,
$\ddot{D}+2H\dot{D}-4\pi G\bar{\rho}_{m}D=0.$ (75)
During matter domination, we know that the scale factor evolves in time as $a\propto t^{2/3}$ and thus $H\equiv\dot{a}/a=\left(2/3\right)t^{-1/3}$. We also know, from the Friedmann equation, that
$4\pi G\bar{\rho}_{m}=\frac{3}{2}H^{2}\Omega_{m}\approx\frac{2}{3}t^{-2}.$ (76)
where we assumed for simplicity $\Omega_{m}=1$. Thus the equation of motion for the SIGF becomes
$\ddot{D}+\frac{4}{3}t^{-1}\dot{D}-\frac{2}{3}t^{-2}D=0.$ (77)
To solve this homogeneous differential equation, we guess $D\left(t\right)\propto t^{n}$ and plug our guess back in Eq. (77),
$n\left(n-1\right)+\frac{4}{3}n-\frac{2}{3}=0,$ (78)
to find that $n$ is either $-1$ or $2/3$. Hence, the SIGF has the following solution
$D\left(t\right)=At^{2/3}+Bt^{-1}\approx At^{2/3},$ (79)
where $A$ and $B$ are constants, to be determined by satisfying the initial conditions for $\delta_{i}$ and $\mathbf{v}_{i}$ at $z=z_{0}$. In the rest of the derivation, we ignore the decaying mode of the SIGF, and assume $D\propto t^{2/3}$.
Next, we solve for $E$, which obeys Eq. (72). During matter domination, this equation becomes
$\ddot{E}+\frac{4}{3}t^{-1}\dot{E}-\frac{2}{3}t^{-2}E=-\frac{2}{3}A^{2}t^{-2/3}.$ (80)
Due to zero initial conditions for $S^{(2)}$, the homogeneous solution for this equation is zero. We guess the following particular solution $E\left(t\right)=cA^{2}t^{m}$. This brings us to
$m\left(m-1\right)ct^{m-2}+\frac{4}{3}mct^{m-2}-\frac{2}{3}ct^{m-2}=-\frac{2}{3}t^{-2/3}.$ (81)
For this equation to hold at any time $t$, we see that we must have $m=4/3$, and thus $E=cA^{2}t^{4/3}=cD^{2}$. In order to find the constant $c$, we plug $m=4/3$ in Eq. (81) and find $c=-3/7$. Therefore, we conclude
$E\left(z\right)=-\frac{3}{7}D^{2}\left(z\right).$ (82)
Note that by relaxing the assumption of $\Omega_{m}=1$, we will get a different $n$ in Eq. (78), which would lead to a different $m$ in Eq. (81) and consequently a different $c$. However, Ref. Bouchet et al. (1995) showed that for a flat universe
$E\left(z\right)\approx-\frac{3}{7}\Omega_{m}^{-1/143}D^{2}\left(z\right)\approx-\frac{3}{7}D^{2}\left(z\right),$ (83)
where the first equality holds to a percent level.
Figure 14: Left panel: realization of the baryon over-density field at $z=20$, evolved non-linearly with scale-dependent 2LPT. Right panel: the fractional error in the normalized density field ($\Delta_{i}\equiv 1+\delta_{i}$) due to approximating $\delta_{b}\approx\delta_{c}$.
Figure 15: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$ (solid curves), $k=0.3\,\mathrm{Mpc}^{-1}$ (dashed curves), and $k=0.15\,\mathrm{Mpc}^{-1}$ (dotted curves), when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). For fair comparison, in all curves we used the non-linear $\delta_{m}^{\mathrm{NL}}$ in the EPS formalism, as in standard 21cmFAST.
D.3 The effect of scale-dependent growth on the 21-cm signal at cosmic dawn
In 21cmFAST, the SIGF approximation of $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$ is taken, and Eqs. (62), (74) are implemented in the code for computing the first and second order expressions for the displacement vector field. In order to compute the non-linear baryon density field more precisely, we replaced Eq. (62) in 21cmFirstCLASS with Eq. (61), i.e. we compute the first order expression for $\mathbf{s}_{i}^{(1)}$ via the SDGF $\mathcal{D}_{i}\left(\mathbf{k},z\right)$, for each species $i$. As for the second order expression, $\mathbf{s}_{i}^{(2)}$, we did not attempt to solve the coupled integro-differential equations given by Eq. (D.2), and instead we keep using the approximated Eq. (74). This is justified as by construction the second order correction $\mathbf{s}_{i}^{(2)}$ has a smaller contribution to $\mathbf{s}_{i}$, compared to the first order correction $\mathbf{s}_{i}^{(1)}$. We leave for future work the comparison of the approximated analytical solution of Eq. (74) with the fully numerical solution of Eq. (D.2).
We show in Fig. 14 the comparison between the more precise $\delta_{b}^{\mathrm{NL}}$, computed with the scale-dependent first order LPT, and the common calculation that evolves the density field with the SIGF. At $z=20$, we see that the error is of the order of a percent at the most over- and under-dense regions. At higher redshifts, this error is expected to increase at higher redshifts due to the departure of the SDGF from the SIGF, while at lower redshifts it vanishes as $\mathcal{D}_{b}\left(k,z\right)\to D\left(z\right)$, see Fig. 3.
The error in the evaluation of $\delta_{b}^{\mathrm{NL}}$ with the SIGF is enhanced when the 21-cm power spectrum is considered, due to the intricate interplay between non-linear fluctuations in different fields, e.g. $T_{k}$ and $J_{\alpha}$. Indeed, we see in Fig. 15 that the error in $\Delta_{21}^{2}$ may reach to $\mathcal{O}\left(10\%\right)$ at $z\gtrsim 12$. At low redshifts ($z\lesssim 10$), the error due to this approximation goes to zero because at these redshift $\delta_{b}$ approaches more closely to $\delta_{c}$.

### v2

@@
 The cosmological 21-cm signal is sourced from hyperfine transitions in neutral hydrogen atoms.
-Yet, although hydrogen atoms belong to the sub-class of baryonic matter, it is customary to simplify their treatment and analyze the 21-cm signal as if all the matter in the Universe was in the form of collisionless cold dark matter (CDM).
+Yet, although the abundance of hydrogen atoms follows the baryon density field, semi-numerical codes that simulate the 21-cm signal simplify their treatment as if all the matter in the Universe was in the form of collisionless cold dark matter (CDM).
 This is usually done by evolving the density field via a scale-independent growth factor (SIGF).
@@
 (2014); Ghara and Mellema (2020); Greig et al.
-(2021); Ghara et al.
+(2021); Mittal and Kulkarni (2021); Ghara et al.
 (2021); Maity and Choudhury (2022); Abdurashidova et al.
@@
 (2024); Adi et al.
+(2024); Libanore et al.
 (2025), and possibly shed some light on the nature of dark matter Tashiro et al.
@@
 (2021); Halder et al.
+(2021); Mittal et al.
 (2022); Sarkar et al.
@@
 (2023); Fialkov et al.
-(2023).
+(2023); Mondal and Barkana (2023).
 The growing interest in the 21-cm signal and its analysis motivated the development of a plethora of dedicated codes, designed to estimate the signal and its spatial fluctuations as a function of redshift.
@@
 For example, CAMB Lewis and Challinor (2007) and PowerSpectrum21cm Muñoz et al.
-(2015b) exploit the linearity of the perturbations in this “dark ages” epoch.
+(2015b) exploit the linearity of the perturbations in this “dark ages” epoch, see also Ref.
+Naoz and Barkana (2005).
 After cosmic dawn, once galaxies and stars have formed, the fluctuations in the signal are amplified considerably, and the validity of linear perturbation theory breaks.
@@
 (2003)).
-Figure 1: The global 21-cm signal under different assumptions.
-Green lines are obtained using 21cmFirstCLASS and separate the baryons and dark matter evolution; other lines assume $\delta_{b}\equiv\delta_{c}$.
-In the blue curve, we ran 21cmFAST under the default settings of non-linear evolution; the orange curve instead is obtained setting the density field to evolve linearly, similarly as in Zeus21 (the two codes well agree in these settings).
-Solid (dashed) lines correspond to codes that adopt the non-linear (linear) matter density field in their calculations.
-In both green curves of 21cmFirstCLASS, the baryon density field is evolved non-linearly.
-Figure 2: Comparison of the 21-cm power spectrum on different scales, under the same conditions as in Fig.
-1.
-Here, we show with error bars the expected noise from HERA in its design sensitivity at low redshifts ($z<10$) assuming the moderate foreground scenario, as evaluated by the 21cmSense code Pober et al.
-(2013, 2014) (we used the same 21cmSense settings as detailed in Ref.
-Flitter and Kovetz (2024a)).
-We also show a separate error bar at $z=14$, corresponding to a sensitivity estimate for SKA Meriot and Semelin (2024).
 Semi-numerical codes, whose runtime is shorter by orders of magnitude, are also found in the literature.
@@
 (2011); Muñoz et al.
-(2022) is perhaps the most popular one, having also been used to analyse HERA Phase I data Abdurashidova et al.
-(2022, 2023), but see also simfast21 Santos et al.
+(2022) is perhaps the most popular one, but see also simfast21 Santos et al.
 (2010), CIFOG Hutter (2018) and 21cmSPACE Visbal et al.
@@
 [^1] ARES Mirocha (2014); Mirocha et al.
-(2017) is another speedy code, but unlike Zeus21 and HMreio, it only computes the global 21-cm signal.
+(2017) and ECHO21 Mittal (2025) are also speedy codes, but unlike Zeus21 and HMreio, they only compute the global 21-cm signal.
 Another analytic code, named HMreio Schneider et al.
@@
 To do so, we rely on the newly developed tool 21cmFirstCLASS Flitter and Kovetz (2024a, b).
-This code interfaces CLASS Blas et al.
-(2011) and 21cmFAST, allowing the simulation to begin from recombination with initial conditions that are consistent with the simulated cosmology.
-We find that at the dark ages the approximation $\delta_{b}\approx\delta_{c}$ completely fails, and $\delta_{b}$ must be evolved in a scale-dependent manner in that epoch.
+We find that at the dark ages the approximation $\delta_{b}\approx\delta_{c}$ fails when high order statistics (e.g.
+the power spectrum) are considered, and $\delta_{b}$ must be evolved in a scale-dependent manner in that epoch.
 We also demonstrate, using scale-dependent Lagrangian perturbation theory (LPT), the need to go beyond linear perturbation theory for the baryon over-density field at low redshifts, after cosmic dawn was set.
-Finally, we present how the signal is modified when considering either the “linear” or “non-linear” matter over-density $\delta_{m}$ in the context of the extended Press-Schechter formalism, and we discuss the role of the relative velocity between baryons and dark matter in determining the 21-cm signal.
-Our main results are shown in Figs.
-1-2.
-In Fig.
-2, the $z\lesssim 10$ error bars indicate the expected noise level from HERA in its design sensitivity under the conservative assumption of a moderate foreground scenario, evaluated with 21cmSense Pober et al.
-(2013, 2014), while the $z=14$ error bar illustrates the sensitivity expected for SKA, estimated by Ref.
-Meriot and Semelin (2024).
-This figure demonstrates that the effects discussed in this work should matter in future analyses of 21-cm signal measurements.
+Finally, we present how the signal is modified when considering either the “linear” or “non-linear” matter over-density $\delta_{m}$ in the context of the extended Press-Schechter formalism.
+The modifications made in 21cmFirstCLASS for this work now make it the only public code in the literature that can consistently simulate the fluctuations in the 21-cm signal at all cosmological epochs, from dark ages until reionization, while simultaneously studying the cosmic microwave background (CMB) anisotropies.
 This paper is organized as follows.
+Sec.
+II provides the theoretical background for the 21-cm signal and the matter fields that govern its shape.
 In Sec.
-II, we examine the impact of using the baryon density in the computation of the 21-cm signal during the dark ages.
-We then move to analyzing the role of the baryons and CDM during cosmic dawn and reionization in Sec.
-III.
-We discuss in Sec.
-IV our results from the previous section and write our conclusions in Sec.
-V.
+III we give a short description of the main tool used in our analysis, 21cmFirstCLASS, and the modifications made for this work.
+We show our results in Sec.
+IV, discuss their significance in Sec.
+V, and conclude in Sec.
+VI.
 In addition, this paper contains several appendices.
-In Appendix A, we study the sensitivity of the 21-cm signal to the precise evolution of CDM and the total matter density fields.
-In Appendix B, we study the Boltzmann equation for the spin temperature.
-In Appendix C, we verify the output of 21cmFirstCLASS, by comparing the temperature fluctuations at the dark ages with the analytical prediction from linear perturbation theory.
-Finally, we derive the equations for second-order scale-dependent LPT in Appendix D.
+In Appendix A, we study the Boltzmann equation for the spin temperature.
+In Appendix B, we study the sensitivity of the 21-cm signal to the precise evolution of CDM and the total matter density fields.
+We derive the equations for second-order scale-dependent LPT in Appendix C.
+Finally, in Appendix D, we verify the output of 21cmFirstCLASS, by comparing the temperature fluctuations at the dark ages with the analytical prediction from linear perturbation theory.
 Throughout this work, we adopt cosmological parameters from Planck18 best-fit values Aghanim et al.
@@
 position, redshift, wavenumber, etc) and they should be inferred from the context.
-II Dark ages
-II.1 Matter density fields at the dark ages
-According to the standard cosmological model, two kinds of matter controlled the evolution of the Universe at early times, before the first atoms were formed.
-The first kind are baryons, which include hydrogen and helium nuclei and electrons and protons that interact strongly with photons due to their electric charge.
-The second kind is collisionless CDM.
-While the baryons were subject to radiation pressure on small scales, CDM was able to freely cluster under the influence of self-gravitational forces.
-As the Universe expanded, it eventually reached a low enough temperature such that atoms could have survived immediate ionization for the first time.
-This promptly led to a dramatic increase in the photon mean free path, thus releasing the freely propagating cosmic microwave background (CMB) radiation.
-This key moment in the history of the Universe, dated at redshift $z\approx 1100$, is known as “recombination”.
-Figure 3: Left panel: the scale-dependent growth factor for baryons; note the fluctuations at very high $z$, where the baryon fluid oscillates being still partially coupled to photons.
-Right panel: comparison with the scale-independent growth factor.
-During the dark ages, the fluctuations in the matter density field were still much smaller than unity and linear perturbation theory can be applied in that epoch.
-After recombination took place, on sub-horizon scales larger than the Jeans scale, the linearized continuity and Euler equations for the baryon and CDM fluids are identical Baumann (2022); Dodelson (2003),
-$\dot{\delta}_{b}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{b}$ $=$ 00 0 (1)
-$\dot{\mathbf{v}}_{b}+H\mathbf{v}_{b}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ 00 0 (2)
-$\dot{\delta}_{c}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{c}$ $=$ 00 0 (3)
-$\dot{\mathbf{v}}_{c}+H\mathbf{v}_{c}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ $0,$ (4)
-where $\delta_{b}$ ($\delta_{c}$) is the baryon (CDM) energy over-density, $\mathbf{v}_{b}$ ($\mathbf{v}_{c}$) is the proper baryon (CDM) peculiar velocity, $a$ is the scale factor, and $H\equiv\dot{a}/a$ is the Hubble parameter.
-Overdots denote derivatives with respect to the cosmological time $t$.
-Eqs.
-(1)-(4) are accompanied with the Poisson equation for the gravitational potential $\Phi$,
-$\nabla^{2}\Phi=4\pi G\bar{\rho}_{m}a^{2}\delta_{m},$ (5)
-where $G$ is Newton’s gravitational constant, $\bar{\rho}_{m}$ is the mean matter energy density and $\delta_{m}$ the matter energy over-density,
-$\delta_{m}=\Omega_{m}^{-1}\left(\Omega_{c}\delta_{c}+\Omega_{b}\delta_{b}\right).$ (6)
-We can combine Eqs. (1)-(6) into two coupled second-order differential equations of the same form ($i=b,c$ for baryons and CDM),
-$\ddot{\delta}_{i}+2H\dot{\delta}_{i}-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}=0.$ (7)
-Even though the equations of motion for $\delta_{b}$ and $\delta_{c}$ are the same (at the relevant scales), their solutions are different because of the initial conditions at the time of recombination; prior that moment, CDM has already clustered on all scales, while the energy density of the baryons was oscillating due to the strong coupling with CMB photons.
-Thus, $\delta_{b}\leq\delta_{c}$ on all scales.
-If $\delta_{b}=\delta_{c}=\delta_{m}$, then Eq.
-(7) becomes a single second order differential equation for $\delta_{m}$.
-Since this differential equation is the same on all scales, variable separation can be performed, $\delta_{m}\left(\mathbf{x},z\right)=\delta_{0}\left(\mathbf{x}\right)D\left(z\right)$, where $\delta_{0}\left(\mathbf{x}\right)$ is a function of only the comoving coordinate $\mathbf{x}$ in real space, and $D\left(z\right)$ is the scale-independent growth factor (SIGF), a function of redshift only that obeys the same differential equation for $\delta_{m}$,
-$\ddot{D}+2H\dot{D}-4\pi G\bar{\rho}_{m}D=0,$ (8)
-with the boundary conditions $D\left(z=0\right)=1$ and $D\left(z\to\infty\right)=0$.
-Because $\delta_{c}\geq\delta_{b}$ and $\Omega_{c}>\Omega_{b}$, it is often approximated $\delta_{m}\approx\delta_{c}$ and $\delta_{c}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)D\left(z\right)$ in Fourier space.
-In this work we use a more precise expression for the baryon and CDM over-densities, [^3] Here we assume that $\delta_{b}\approx\delta_{c}=\delta_{0}$ at $z=0$, which is very well justified as at this redshift the baryons have already clustered to the gravitational potential wells sourced by CDM.
-$\delta_{i}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)\mathcal{D}_{i}\left(k,z\right)\qquad i=b,c.$ (9)
-where $\mathcal{D}_{i}\left(k,z\right)$ is the scale-dependent growth factor (SDGF) for species $i$. By construction, it is defined as
-$\mathcal{D}_{i}\left(k,z\right)\equiv\frac{\mathcal{T}_{i}\left(k,z\right)}{\mathcal{T}_{i}\left(k,z=0\right)}\qquad i=b,c.$ (10)
-where $\mathcal{T}_{i}\left(k,z\right)=\delta_{i}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the matter density transfer function of species $i$ and $\mathcal{R}$ is the primordial curvature perturbation.
-The CDM SDGF is nearly scale-invariant and can be very well approximated by $\mathcal{D}_{c}\left(k,z\right)\approx D\left(z\right)$, apart for BAO features that appear at $z\lesssim 1$ (while adopting the SIGF for the CDM density field causes “fake BAOs” at high redshifts, see more details in Appendix A).
-In contrast, the baryon SDGF $\mathcal{D}_{b}\left(k,z\right)$ is much smaller than $D\left(z\right)$ at high redshifts, while at low redshifts they are nearly the same due to the increasing coupling due to gravitation between the baryon and CDM fluids.
-This is demonstrated in Fig.
-3, where BAO features are clearly manifest.
-Since $\mathcal{D}_{b}\left(k,z\right)\ll\mathcal{D}_{c}\left(k,z\right)$ at the dark ages, we have $\delta_{b}\left(\mathbf{x},z\right)\ll\delta_{c}\left(\mathbf{x},z\right)$ from Eq.
-(9), as illustrated in Fig 4.
-This has some important implications on the 21-cm signal at that epoch.
-Figure 4: Left (right) panel: realization of the CDM (baryon) over-density field at $z=300$.
-II.2 Brightness temperature at the dark ages
+II Theory
+In this section, we summarize the main theoretical components needed to consistently model the evolution of the 21-cm signal during the dark ages and cosmic dawn.
+All of them, have been implemented in the public code 21cmFirstCLASS, as described in Sec.
+III.
+II.1 The 21-cm signal
 The quantity of most interest in the study of the 21-cm signal is the brightness temperature relative to the CMB, defined as
@@
 In theory, the solution for $T_{s}$ has to be obtained by solving the Boltzmann equation for $n_{1}$ and $n_{0}$.
-Yet, we show in Appendix B that the solution for this equation is very well approximated by the well known expression for $T_{s}$ in equilibrium, even in the dark ages where $H\gg A_{10}$.
+Yet, we show in Appendix A that the solution for this equation is very well approximated by the well known expression for $T_{s}$ in equilibrium, even in the dark ages where $H\gg A_{10}$.
 This equilibrium expression reads
@@
 $+\left.\frac{dT_{k}}{dt}\right|_{\mathrm{CMB}}+\left.\frac{dT_{k}}{dt}\right|_{X},$
-where $x_{e}\equiv n_{e}/\left(n_{\mathrm{H}}+n_{\mathrm{He}}\right)$ is the free-electron fraction, with $n_{e}$, $n_{\mathrm{H}}$ and $n_{\mathrm{He}}$ the free electron, hydrogen nuclei and helium nuclei number densities, respectively.
+where $\delta_{b}$ is the baryon energy over-density and $x_{e}\equiv n_{e}/\left(n_{\mathrm{H}}+n_{\mathrm{He}}\right)$ is the free-electron fraction, with $n_{e}$, $n_{\mathrm{H}}$ and $n_{\mathrm{He}}$ the free electron, hydrogen nuclei and helium nuclei number densities, respectively.
 The first line in Eq.
@@
 21cmFirstCLASS involves the state of the art code HYREC Ali-Haimoud and Hirata (2011); Lee and Ali-Haïmoud (2020) that solves with great precision the fully-detailed Boltzmann equation for $x_{e}$ while accounting for helium recombination and high-order energy excitations.
-During the dark ages, the only relevant matter species that contributes to the 21-cm signal are the baryons, as they source the emission and absorption of 21-cm radiation by hydrogen atoms, the recombination rate that governs $T_{k}$, and the coupling of $T_{s}$ to $T_{k}$ via collisional excitations.
-Thus the fluctuations in the 21-cm signal during the dark ages are entirely proportional to $\delta_{b}$.
-Because $\delta_{b}\ll 1$ during the dark ages, the scale-dependent evolution of the baryon density field barely affects the global 21-cm signal, i.e.
-the sky-averaged brightness temperature as a function of redshift.
+II.2 Growth factor
+According to the standard cosmological model, two kinds of matter controlled the evolution of the Universe at early times, before the first atoms were formed.
+The first kind are baryons, which include hydrogen and helium nuclei and electrons [^3] We follow standard nomenclature and cavalierly group electrons with baryons.
+and protons that interact strongly with photons due to their electric charge.
+The second kind is collisionless CDM.
+While the baryons were subject to radiation pressure on small scales, CDM was able to freely cluster under the influence of self-gravitational forces.
+As the Universe expanded, it eventually reached a low enough temperature such that neutral atoms formed for the first time.
+This promptly led to a dramatic increase in the photon mean free path, thus releasing the freely propagating CMB radiation.
+This key moment in the history of the Universe, dated at redshift $z\approx 1100$, is known as “recombination”.
+Figure 1: Left panel: the scale-dependent growth factor for baryons; note the fluctuations at very high $z$, where the baryon fluid oscillates being still partially coupled to photons.
+Right panel: comparison with the scale-independent growth factor.
+During the dark ages, the fluctuations in the matter density field were still much smaller than unity and linear perturbation theory can be applied in that epoch.
+After recombination took place, the linearized continuity and Euler equations for the baryon and CDM fluids on sub-horizon scales are Baumann (2022); Dodelson (2003),
+$\dot{\delta}_{b}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{b}$ $=$ 00 0 (5)
+$\dot{\mathbf{v}}_{b}+H\mathbf{v}_{b}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ $-a^{-1}c_{s}^{2}\boldsymbol{\nabla}\delta_{b}$ (6)
+$\dot{\delta}_{c}+a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{c}$ $=$ 00 0 (7)
+$\dot{\mathbf{v}}_{c}+H\mathbf{v}_{c}+a^{-1}\boldsymbol{\nabla}\Phi$ $=$ $0,$ (8)
+where $\delta_{c}$ is the CDM energy over-density, $\mathbf{v}_{c}$ is the proper CDM peculiar velocity, and $c_{s}$ the baryonic sound speed.
+Overdots denote derivatives with respect to the cosmological time $t$.
+Eqs.
+(5)-(8) are accompanied with the Poisson equation for the gravitational potential $\Phi$,
+$\nabla^{2}\Phi=4\pi G\bar{\rho}_{m}a^{2}\delta_{m},$ (9)
+where $G$ is Newton’s gravitational constant, $\bar{\rho}_{m}$ is the mean matter energy density and $\delta_{m}$ the matter energy over-density,
+$\delta_{m}=\Omega_{m}^{-1}\left(\Omega_{c}\delta_{c}+\Omega_{b}\delta_{b}\right).$ (10)
+Below the comoving Jeans scale, $\sim 10\,\mathrm{kpc}$, the fluctuations of the 21-cm signal are suppressed due to smoothing caused by the pressure of the baryonic gas Loeb and Zaldarriaga (2004); Naoz and Barkana (2005), as is manifest through the $c_{s}$ term in Eq.
+(6).
+We have showed in a previous work Flitter and Kovetz (2024b) that above the Jeans scale, at a wavenumber $k\lesssim 50\,\mathrm{Mpc}^{-1}$, the fluctuations in the 21-cm signal can be very well captured by omitting this term.
+Under this approximation, the equations of motion for the baryon and CDM fluids become identical and we can combine Eqs.
+(5)-(8) into two coupled second-order differential equations of the same form ($i=b,c$ for baryons and CDM),
+$\ddot{\delta}_{i}+2H\dot{\delta}_{i}-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}\delta_{j}=0.$ (11)
+Even though the equations of motion for $\delta_{b}$ and $\delta_{c}$ are the same (at the relevant scales), their solutions are different because of the initial conditions at the time of recombination; prior that moment, CDM has already clustered on all scales, while the energy density of the baryons was oscillating due to the strong coupling with CMB photons.
+Thus, $\delta_{b}\leq\delta_{c}$ on all scales.
+If $\delta_{b}=\delta_{c}=\delta_{m}$, then Eq.
+(11) becomes a single second order differential equation for $\delta_{m}$.
+Since this differential equation is the same on all scales, variable separation can be performed, $\delta_{m}\left(\mathbf{x},z\right)=\delta_{0}\left(\mathbf{x}\right)D\left(z\right)$, where $\delta_{0}\left(\mathbf{x}\right)$ is a function of only the comoving coordinate $\mathbf{x}$ in real space, and $D\left(z\right)$ is the scale-independent growth factor (SIGF), a function of redshift only that obeys the same differential equation for $\delta_{m}$,
+$\ddot{D}+2H\dot{D}-4\pi G\bar{\rho}_{m}D=0,$ (12)
+with the boundary conditions $D\left(z=0\right)=1$ and $D\left(z\to\infty\right)=0$.
+Because $\delta_{c}\geq\delta_{b}$ and $\Omega_{c}>\Omega_{b}$, it is often approximated $\delta_{m}\approx\delta_{c}$ and $\delta_{c}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)D\left(z\right)$ in Fourier space.
+In this work we use a more precise expression for the baryon and CDM over-densities, [^4] Here we assume that $\delta_{b}\approx\delta_{c}=\delta_{0}$ at $z=0$, which is very well justified as at this redshift the baryons have already clustered to the gravitational potential wells sourced by CDM.
+$\delta_{i}\left(\mathbf{k},z\right)=\delta_{0}\left(\mathbf{k}\right)\mathcal{D}_{i}\left(k,z\right)\qquad i=b,c.$ (13)
+where $\mathcal{D}_{i}\left(k,z\right)$ is the scale-dependent growth factor (SDGF) for species $i$. By construction, it is defined as
+$\mathcal{D}_{i}\left(k,z\right)\equiv\frac{\mathcal{T}_{i}\left(k,z\right)}{\mathcal{T}_{i}\left(k,z=0\right)}\qquad i=b,c.$ (14)
+where $\mathcal{T}_{i}\left(k,z\right)=\delta_{i}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the matter density transfer function of species $i$ and $\mathcal{R}$ is the primordial curvature perturbation.
+The CDM SDGF is nearly scale-invariant and can be very well approximated by $\mathcal{D}_{c}\left(k,z\right)\approx D\left(z\right)$, apart for BAO features that appear at $z\lesssim 1$ (while adopting the SIGF for the CDM density field causes “fake BAOs” at high redshifts, see more details in Appendix B).
+In contrast, the baryon SDGF $\mathcal{D}_{b}\left(k,z\right)$ is much smaller than $D\left(z\right)$ at high redshifts, while at low redshifts they are nearly the same due to the increasing coupling due to gravitation between the baryon and CDM fluids.
 This is demonstrated in Fig.
-5, where the baryon over-density field is evolved with the SIGF (blue curve), or the SDGF (orange curve).
-Figure 5: The 21-cm global signal as a function of redshift, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
-The curves almost completely overlap each other.
-Although the scale-dependent evolution of $\delta_{b}$ does not affect the global 21-cm signal, it does affect its two-point statistics. Specifically, let us consider the 21-cm power spectrum,
-$\Delta_{21}^{2}\left(k,z\right)=\frac{k^{3}\bar{T}_{21}^{2}\left(z\right)P_{21}\left(k,z\right)}{2\pi^{2}},$ (15)
-where $\bar{T}_{21}$ is the global brightness temperature and $P_{21}\left(k,z\right)$ is the angle-averaged Fourier transform of the two-point correlation function $\langle\delta_{21}\left(\mathbf{x},z\right)\delta_{21}\left(\mathbf{x}^{\prime},z\right)\rangle$, while $\delta_{21}$ is the local contrast in the brightness temperature, $\delta_{21}\left(\mathbf{x},z\right)\equiv T_{21}\left(\mathbf{x},z\right)/\bar{T}_{21}\left(z\right)-1$.
-In a previous work Flitter and Kovetz (2024b), we have shown that during the dark ages $\Delta_{21}^{2}\left(k,z\right)\propto\mathcal{T}_{b}^{2}\left(k,z\right)\ll\mathcal{T}_{c}^{2}\left(k,z\right)$.
-The analysis in that work was based on the same SDGF formalism we present here, which we have shown to agree well with fully relativistic Boltzmann-Einstein solvers.
-In Fig.
-6, we show $\Delta_{21}^{2}\left(k,z\right)$, as calculated by 21cmFirstCLASS after the incorporation of the SDGF in the code (we verify the correctness of our results in Appendix C).
-As expected, $\Delta_{21}^{2}\left(k,z\right)$ is suppressed during the dark ages when $\delta_{b}$ is correctly evolved with the SDGF.
-Figure 6: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
-For a zoomed-in version of the figure at cosmic dawn, see Fig.
-15 in Appendix D.
-III Cosmic dawn and reionization
-III.1 Baryons
-The role of the baryons in shaping the 21-cm signal is enhanced when the dark ages are over.
-As dark matter halos collapse, the gravitational pull that the baryons experience is increased.
-If the cooling timescale is shorter than the free-fall timescale, zero-metalicity popIII stars can form via gas accretion, mostly in molecular cooling galaxies (MCGs) Tegmark et al.
-(1997); Abel et al.
-(2002); Bromm and Larson (2004); Haiman and Bryan (2006); Trenti (2010); Plombat and Puy (2024).
-Baryonic feedbacks, such as supernovae Springel and Hernquist (2003); Wise and Abel (2008); Pawlik and Schaye (2009), photo-heating Hui and Gnedin (1997); Okamoto et al.
-(2008); Ocvirk et al.
-(2020); Katz et al.
-(2020) and Lyman-Werner radiation Bromm and Larson (2004); Fialkov et al.
-(2013); Haiman et al.
-(1997) hamper the star formation rate (SFR).
-Still, stars continue to form in more massive halos that host atomic cooling galaxies (ACGs) Qin et al.
-(2020, 2021a); Krumholz (2012).
-Those halos typically are assumed to contain mainly popII stars with low metalicity.
-[…]
-The local SFRD does not depend only on the local baryon density, but also on local velocities, specifically on the relative velocity between baryons and CDM, $\mathbf{v}_{cb}=\mathbf{v}_{c}-\mathbf{v}_{b}$.
-From Eqs.
-(2) and (4), we see that $\dot{\mathbf{v}}_{cb}=-H\mathbf{v}_{cb}$, which has the temporal solution $\mathbf{v}_{cb}\left(\mathbf{x},z\right)\propto\left(1+z\right)$.
-As for the spatial solution, adiabatic initial conditions imply Gaussian distributions for each one of the three components of $\mathbf{v}_{cb}$, and a Maxwell-Boltzmann distribution for $v_{cb}\equiv\left(\mathbf{v}_{cb}\cdot\mathbf{v}_{cb}\right)^{1/2}$.
-The variance of this distribution can be determined by defining $\theta_{cb}\equiv a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{cb}$, and then
-$\langle v_{cb}^{2}\left(z\right)\rangle=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\frac{\mathcal{T}^{2}_{\theta_{cb}}\left(k,z\right)}{k^{2}},$ (16)
-[…]
-Figure 7: The effect of the baryon-CDM relative velocity $v_{cb}$ on the 21-cm global signal.
-In the green curve, the effect of $v_{\mathrm{cb}}$ enters only through the modification of the minimum halo mass that can host popIII stars, as implemented in 21cmFirstCLASS.
-In the orange curve, small-scale suppression effects in $\sigma\left(M_{h},z\right)$ are also included, as in 21cmFAST, see more details in Sec.
-IV.
-The small extra delay between the orange and green curves can be seen more clearly in Fig.
-1 (as the difference between the solid and dashed green curves at $z\gtrsim$10).
+1, where BAO features are clearly manifest.
+Since $\mathcal{D}_{b}\left(k,z\right)\ll\mathcal{D}_{c}\left(k,z\right)$ at the dark ages, we have $\delta_{b}\left(\mathbf{x},z\right)\ll\delta_{c}\left(\mathbf{x},z\right)$ from Eq.
+(13), as illustrated in Fig 2.
+This has some important implications on the 21-cm signal at that epoch, as we show in Sec IV.1.
+Figure 2: Left (right) panel: realization of the CDM (baryon) over-density field at $z=300$. The color bar is the same for the two plots, hence highlighting the stronger fluctuations in the CDM field.
+II.3 Non linear evolution
 Once galaxies have started to form inside halos, the baryon density field is far from homogeneous as $\delta_{b}>1$ in regions of galaxy clusters.
@@
 (1995); Buchert (1995); Tatekawa (2005); Matsubara (2008); Carlson et al.
-(2009); Zheligovsky and Frisch (2014) (2LPT), as in 21cmFAST, though we improved its calculations in 21cmFirstCLASS to account for the scale-dependent growth of the baryons.
-Readers who wish to learn more about the scale-dependent 2LPT can find more details in Appendix D.
-The importance of evolving the baryon density field non-linearly is evident in Fig.
-8; while linear perturbation theory seems to give a rough estimate for the distribution of baryonic matter on very large scales, it fails to capture the clustering of baryons at low redshifts and small scales, and it wrongly predicts a more homogeneous density field.
-When computing the 21-cm power spectrum, the imprecise calculation from linear perturbation theory yields an order unity error at $z\lesssim 8$, as demonstrated in Fig.
-9.
-Importantly, this is the regime where upcoming 21-cm interferometers are mostly sensitive (c.f.
-Fig.
-2).
-Figure 8: Left (right) panel: realization of the baryon over-density field at $z=6$ when evolved linearly (non-linearly).
-Figure 9: The 21-cm power spectrum as a function of redshift at $k=0.3\,\mathrm{Mpc}^{-1}$, when the baryon and the total matter over-density fields are evolved either linearly or non-linearly.
-The colors of the curves here match the same colors in Figs.
-1-2, namely the blue, orange and green curves can be associated with 21cmFAST, Zeus21-like, and 21cmFirstCLASS, respectively.
-Here though, unlike Figs.
-1-2, in all curves we distinguish between baryons and CDM (i.e.
-in all curves we use the SDGF), and we have the same $v_{cb}$ delaying effect (see differences in Fig.
-7).
-III.2 Total matter
+(2009); Zheligovsky and Frisch (2014) (2LPT).
+We provide in Appendix C the full derivation of the scale-dependent 2LPT equations.
+In brief, in this method parcels move from their initial Lagrangian coordinates $\mathbf{q}$ at redshift $z_{0}$ to their Eulerian coordinates $\mathbf{x}$ according the displacement vector field $\mathbf{s}$, such that $\mathbf{x}\left(\mathbf{q},z\right)=\mathbf{q}\left(z_{0}\right)+\mathbf{s}\left(\mathbf{q},z\right)$.
+The solution to $\mathbf{s}$ can then be expanded in perturbation theory, $\mathbf{s}=\mathbf{s}^{(1)}+\mathbf{s}^{(2)}$.
+Under the SIGF approximation, in which $\delta_{b}=\delta_{c}$, $\mathbf{s}^{(1)}$ and $\mathbf{s}^{(2)}$ are given by
+$\mathbf{s}^{\left(1\right)}\left(\mathbf{q},z\right)$ $=$ $\left[D\left(z\right)-D\left(z_{0}\right)\right]\boldsymbol{\nabla}_{\mathbf{q}}\psi^{\left(1\right)}\left(\mathbf{q}\right)$ (15)
+$\mathbf{s}^{\left(2\right)}\left(\mathbf{q},z\right)$ $=$ $\left[E\left(z\right)-E\left(z_{0}\right)\right]\boldsymbol{\nabla}_{\mathbf{q}}\psi^{\left(2\right)}\left(\mathbf{q}\right),$ (16)
+where
+$E\left(z\right)=-\frac{3}{7}D^{2}\left(z\right),\qquad\nabla_{\mathbf{q}}^{2}\psi^{(1)}\left(\mathbf{q}\right)=-\delta_{0}\left(\mathbf{q}\right),$ (17)
+$\nabla_{\mathbf{q}}^{2}\psi^{\left(2\right)}=\frac{1}{2}\left[\left(\nabla_{\mathbf{q}}^{2}\psi^{\left(1\right)}\right)^{2}-\left(\frac{\partial^{2}\psi^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (18)
+In Sec. III, we discuss how these equations change when the SDGF is introduced.
+II.4 Conditional mass function
 During the cosmic dawn and reionization epochs, the 21-cm signal does not depend only on the baryon over-density $\delta_{b}$, but also on the CDM over-density field $\delta_{c}$, or more precisely on the total matter over-density $\delta_{m}$, Eq.
 (10).
-This is because the SFRD is given by the following expression Madau et al.
-(1997); Schneider et al.
-(2021); Muñoz et al.
-(2022); Muñoz (2023); Cruz et al.
-(2024b),
-[…]
-where $\dot{M}_{*}\left(\mathbf{x},M_{h},z\right)$ is the local stellar mass production rate in a halo of mass $M_{h}$ and $dn\left(\mathbf{x},M_{h},z\right)/dM_{h}$ is the conditional halo mass function (HMF); it is conditional in the sense that it depends on the value of the local total matter over-density $\delta_{m}\left(\mathbf{x},z\right)$.
+This is because galaxies and stars, which emit crucial radiation that alters the signal, form inside collapsed dark matter halos.
+In analytical and semi-numerical models, the effect of the halos on the signal comes into play through the conditional halo mass function (HMF), $dn\left(\mathbf{x},M_{h},z\right)/dM_{h}$; it is conditional in the sense that it depends on the value of the local total matter over-density $\delta_{m}\left(\mathbf{x},z\right)$.
 The mean of the conditional HMF is the “standard” HMF, defined as Lukic et al.
@@
 Perhaps the simplest choice is to rely on the purely Gaussian fitting function from the Press-Schechter (PS) Press and Schechter (1974) HMF.
-In this work, we use instead the Sheth-Tormen (ST) HMF Sheth and Tormen (1999); Sheth et al.
+While the simple form of the PS fitting function captures well the qualitative behavior of the HMF, it under-estimates the number of massive halos.
+For that reason, we use in this work instead the Sheth-Tormen (ST) HMF Sheth and Tormen (1999); Sheth et al.
 (2001); Sheth and Tormen (2002) (but see other popular mass functions in Refs.
@@
 (2013)).
-In evaluating the integral in Eq. (17), the following quantity is often considered,
+The calculation of the emissivity of the radiating sources often involves the conditional halo mass function through the integral in Eq.
+(25).
+In evaluating that integral, the following quantity can be useful,
 […]
 This is the local collapsed fraction of halos more massive than $M_{\mathrm{min}}\left(z\right)$.
+From Eq.
+(21) it is clear then that the conditional HMF can be extracted from the collapsed fraction through the relation
+$\frac{dn\left(\mathbf{x},z,M_{h}\right)}{dM_{h}}=-\frac{\bar{\rho}_{m,0}}{M_{h}}\frac{df_{\mathrm{coll}}\left(\mathbf{x},z;M_{h}\right)}{dM_{h}}.$ (22)
 The collapsed fraction is a very useful quantity since the extended Press-Schechter (EPS) theory predicts its value analytically Bond et al.
@@
 Note that in writing Eq.
-(III.2) we have used the redshift-dependent $\sigma\left(M_{h},z\right)$ as given by Eq.
-(19), without imposing the SIGF approximation, $\sigma\left(M_{h},z\right)\propto D\left(z\right)$, which is less accurate at high redshifts and low masses (see more details in Appendix A).
-IV Discussion
+(II.4) we have used the redshift-dependent $\sigma\left(M_{h},z\right)$ as given by Eq.
+(20), without imposing the SIGF approximation, $\sigma\left(M_{h},z\right)\propto D\left(z\right)$, which is less accurate at high redshifts and low masses (see more details in Appendix B).
 One delicate nuance regarding the formalism presented above concerns the assumed matter density field that is used in Eq.
-(III.2).
+(II.4).
 Press-Schechter theory Sheth and Tormen (1999), and the extended Press-Schechter theory Bond et al.
@@
 Hence, from a purely theoretical point of view, it would be more consistent to evaluate the smoothed $\delta_{m}$ that appears in Eq.
-(III.2), as well as $\sigma\left(M_{h},z\right)$, from linear theory.
+(II.4), as well as $\sigma\left(M_{h},z\right)$, from linear theory.
 This is for example the underlying assumption in Zeus21, where the “linear” $\delta_{m}^{\mathrm{lin}}$ still results in non-linear fluctuations in the SFRD due to the erfc function in Eq.
-(III.2), an effect that can be captured by an exponential fit Muñoz (2023).
+(II.4), an effect that can be captured by an exponential fit Muñoz (2023).
 In contrast, 21cmFAST uses the non-linear $\delta_{m}^{\mathrm{NL}}$ (from 2LPT calculations) in Eq.
-(III.2).
+(II.4).
 The justification for this choice comes from better agreement with radiative transfer simulations at the epoch of reionization Zahn et al.
@@
 (2013).
-Evidently, because the fluctuations in the brightness temperature are strongly connected to the fluctuations in the SFRD, the choice of working with either $\delta_{m}^{\mathrm{lin}}$ or $\delta_{m}^{\mathrm{NL}}$ has large implications on the resulting 21-cm power spectrum, as can be seen in Fig.
-9.
-Our readers are hence forewarned.
-The second nuance that we wish to highlight is the effect that $v_{cb}$ has on the local SFRD.
-As previously discussed, the local SFRD is suppressed in regions with large relative velocity.
-In 21cmFAST and 21cmFirstCLASS this effect is modelled in the code by increasing $M_{\mathrm{min}}$ in MCGs Muñoz et al.
-(2022),
+II.5 Star formation rate density
+The role of the baryons in shaping the 21-cm signal is enhanced when the dark ages are over.
+As dark matter halos collapse, the gravitational pull that the baryons experience is increased.
+If the cooling timescale is shorter than the free-fall timescale, zero-metalicity popIII stars can form via gas accretion, mostly in molecular cooling galaxies (MCGs, other works often refer to molecular cooling halos for more generality Tegmark et al.
+(1997); Abel et al.
+(2002); Bromm and Larson (2004); Haiman and Bryan (2006); Trenti (2010); Plombat and Puy (2024)).
+Baryonic feedbacks, such as supernovae Springel and Hernquist (2003); Wise and Abel (2008); Pawlik and Schaye (2009), photo-heating Hui and Gnedin (1997); Okamoto et al.
+(2008); Ocvirk et al.
+(2020); Katz et al.
+(2020) and Lyman-Werner radiation Bromm and Larson (2004); Fialkov et al.
+(2013); Haiman et al.
+(1997) hamper the star formation rate (SFR).
+Still, stars continue to form in more massive halos that host atomic cooling galaxies (ACGs) Qin et al.
+(2020, 2021a); Krumholz (2012).
+Those halos typically are assumed to contain mainly popII stars with low metallicity, though popIII stars can still form in atomic cooling halos if there is sterilizing radiation background that prevents star formation earlier in the molecular cooling stage (when the halo was smaller and younger).
+[…]
+[…]
+where $\dot{M}_{*}\left(\mathbf{x},M_{h},z\right)$ is the local stellar mass production rate in a halo of mass $M_{h}$.
+Clearly, the local SFRD is smaller in regions with smaller $\dot{M}_{*}$, while the latter quantity must be considerably low at halos of mass $\sim M_{\mathrm{min}}$, where the halo mass is insufficient for efficient cooling.
+In practice, Eqs.
+(22) and (II.4) are used in computing $\dot{\rho}_{*}\left(\mathbf{x},z\right)$ under the assumption of PS HMF, while the unconditional ST HMF is then used to normalize the mean SFRD Barkana and Loeb (2004, 2008); Mesinger et al.
+(2011).
+The local SFRD does not depend only on the local baryon density, but also on local velocities, specifically on the relative velocity between baryons and CDM, $\mathbf{v}_{cb}=\mathbf{v}_{c}-\mathbf{v}_{b}$.
+From Eqs.
+(6) and (8), we see that above the Jeans scale $\dot{\mathbf{v}}_{cb}=-H\mathbf{v}_{cb}$, which has the temporal solution $\mathbf{v}_{cb}\left(\mathbf{x},z\right)\propto\left(1+z\right)$.
+As for the spatial solution, adiabatic initial conditions imply Gaussian distributions for each one of the three components of $\mathbf{v}_{cb}$, and a Maxwell-Boltzmann distribution for $v_{cb}\equiv\left(\mathbf{v}_{cb}\cdot\mathbf{v}_{cb}\right)^{1/2}$.
+The variance of this distribution can be determined by defining $\theta_{cb}\equiv a^{-1}\boldsymbol{\nabla}\cdot\mathbf{v}_{cb}$, and then
+$\langle v_{cb}^{2}\left(z\right)\rangle=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\frac{\mathcal{T}^{2}_{\theta_{cb}}\left(k,z\right)}{k^{2}},$ (26)
+where $\mathcal{T}_{\mathrm{\theta_{cb}}}\left(k,z\right)=\theta_{cb}\left(k,z\right)/\mathcal{R}\left(k\right)$ is the transfer function for $\theta_{cb}$.
+Evaluating this integral at recombination ($z\approx 1100$) reveals that baryons move with a mean relative velocity of $\sim 30\,\mathrm{km/sec}$ with respect to CDM, corresponding to a Mach number of 5 Tseliakhovich and Hirata (2010); Ali-Haïmoud et al.
+(2014).
+This supersonic velocity makes it harder for the baryons to collapse and form popIII stars in small dark matter halos Tseliakhovich et al.
+(2011); Dalal et al.
+(2010); Fialkov et al.
+(2012); Greif et al.
+(2011); Stacy et al.
+(2011); O’Leary and McQuinn (2012); Schauer et al.
+(2019); Hirano et al.
+(2018); Barkana (2016); Muñoz (2019a), thereby lowering the SFRD and delaying cosmic dawn.
+[…]
 $M_{\mathrm{min}}^{\mathrm{MCG}}\propto\left[1+A_{v_{cb}}\frac{v_{cb}\left(\mathbf{x},z_{\mathrm{rec}}\right)}{\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle^{1/2}}\right]^{\beta_{v_{cb}}},$ (27)
 where $A_{v_{cb}}$ and $\beta_{v_{cb}}$ are free parameters, and $\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle$ is the variance of the $v_{cb}$ field at recombination (see Eq.
-(16)).
+26).
 This modification in $M_{\mathrm{min}}$ results in having VAOs in the 21-cm power spectrum at large scales Muñoz (2019a, b); Sarkar and Kovetz (2023); Zhang et al.
+(2024), as was first studied in Refs.
+Dalal et al.
+(2010); Visbal et al.
+(2012); Fialkov et al.
 (2013).
-In addition to the above modification, 21cmFAST also modifies the current matter power spectrum by introducing a multiplicative correction factor of the form $1-A_{p}\exp\left[-\left(\ln\left[k/k_{p}\right]\right)^{2}/\left(2\sigma_{p}^{2}\right)\right]$ Muñoz et al.
+III Methods
+The main tool we use in this work to study how the 21-cm signal is sensitive to the assumed underlying matter fields is 21cmFirstCLASS.
+This code interfaces CLASS Blas et al.
+(2011) and 21cmFAST, allowing the simulation to begin from recombination with initial conditions that are consistent with the simulated cosmology, evolve the matter fields throughout the dark ages, and apply 2LPT during cosmic dawn and reionization to simulate the non-linear fluctuations of the signal in those epochs.
+Moreover, the merger between CLASS and 21cmFAST enables a consistent study of both CMB and 21-cm power spectra with 21cmFirstCLASS, even for cosmologies that extend beyond $\Lambda$CDM Flitter and Kovetz (2024a); Adi et al.
+(2024); Libanore et al.
+(2025).
+For this work, three major modifications have been made in 21cmFirstCLASS with respect to its previous version, whose full description can be found in Ref.
+Flitter and Kovetz (2024a).
+Firstly, we enabled the option of using the linear matter density field $\delta_{m}^{\mathrm{lin}}$, rather the non-linear matter density field $\delta_{m}^{\mathrm{NL}}$ in the evaluation of the inhomogeneous SFRD $\dot{\rho}_{*}\left(\mathbf{x},z\right)$ (see discussion at the end of Sec.
+II.4).
+Secondly, we changed the effect that $v_{cb}$ has on the local SFRD; in addition to the modification of $M_{\mathrm{min}}^{\mathrm{MCG}}$ by $v_{cb}$ (Eq.
+27), 21cmFAST also modifies the current matter power spectrum by introducing a multiplicative correction factor of the form $1-A_{p}\exp\left[-\left(\ln\left[k/k_{p}\right]\right)^{2}/\left(2\sigma_{p}^{2}\right)\right]$ Muñoz et al.
 (2022).
 This factor arises from non-linear $v_{\mathrm{cb}}$ corrections, and was calibrated at $z=20$ with $A_{p}=0.24$, $k_{p}=300\,\mathrm{Mpc}^{-1}$, and $\sigma_{p}=0.9$.
-21cmFirstCLASS has the option to remove this factor since it introduces non-linear corrections to $\delta_{m}^{\mathrm{lin}}$ and $\sigma\left(M_{h},z\right)$, while the EPS formalism requires these quantities to be evaluated from linear theory.
-Because this factor suppresses power at small scales, removing it advances cosmic dawn to slightly higher redshifts, as can be seen in Fig.
-1 and 7.
-A second forewarning due here.
-Finally, a third nuance in the calculation arises due to the appearance of the smoothing radius $R$ in the RHS of Eq.
-(III.2).
+21cmFirstCLASS has the option to remove this factor since it introduces non-linear corrections to $\delta_{m}^{\mathrm{lin}}$ and $\sigma\left(M_{h},z\right)$, while the EPS formalism requires these quantities to be evaluated from linear theory, as we now do in 21cmFirstCLASS.
+Figure 3: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
+The inset shows a zoomed-in version of the curves after the dark ages, where a maximum difference of 6% can be seen.
+Figure 4: Top left panel: realization of the baryon over-density field at $z=20$, evolved non-linearly with scale-dependent 2LPT.
+Top right panel: the fractional error in the normalized density field ($\Delta_{i}\equiv 1+\delta_{i}$) due to approximating $\delta_{b}\approx\delta_{c}$.
+The error is of the order of a percent at the most over- and under-dense regions.
+Bottom left (right) panel: realization of the baryon over-density field at $z=6$ when evolved linearly (non-linearly).
+Linear perturbation theory fails to capture the clustering of baryons at low redshifts and small scales.
+The third modification that was made in 21cmFirstCLASS is relaxing the SIGF approximation in the calculations by accounting for the more exact SDGF.
+Assuming that linear perturbation theory holds during the dark ages, we have shown in Ref.
+Flitter and Kovetz (2024b) that implementing the SDGF in 21cmFirstCLASS would yield excellent agreement with fully relativistic Boltzmann-Einstein solvers such as CAMB, on sub-horizon scales that are larger than the Jeans scale (we verify the correctness of our calculations in Appendix D).
+On top of this modification during the dark ages, we improved the 2LPT calculations in 21cmFAST to account for scale-dependent growth also in the non-linear epochs of cosmic dawn and reionization.
+When relaxing the SIGF approximation, the displacement vector field depends on the considered matter species, and we denote it by $\mathbf{s}_{i}\left(\mathbf{q},z\right)$ ($i=b,c$).
+Now, the first order correction $\mathbf{s}_{i}^{(1)}\left(\mathbf{q},z\right)$ cannot be decomposed to spatial and temporal components, as in Eq.
+(15), and instead we must compute it in Fourier space,
+$\mathbf{s}_{i}^{\left(1\right)}\left(\mathbf{k},z\right)=\left[\mathcal{D}_{i}\left(k,z\right)-\mathcal{D}_{i}\left(k,z_{0}\right)\right]\frac{i\mathbf{k}}{k^{2}}\delta_{0}\left(\mathbf{k}\right).$ (28)
+When we relax the SIGF approximation in 21cmFirstCLASS, we use Eq.
+(28) for $\mathbf{s}_{i}^{(1)}$, instead of Eq.
+(15).
+On the other hand, the second order correction, $\mathbf{s}_{i}^{(2)}$, should be found by solving a set of coupled integro-differential equations, that the interested reader can find in Appendix C, Eq.
+(C.2).
+Instead, in 21cmFirstCLASS we keep using the approximation in Eq.
+(16); as described in detail in Appendix C, this is justified as by construction the second order correction $\mathbf{s}_{i}^{(2)}$ has a smaller contribution to $\mathbf{s}_{i}$, compared to the first order correction $\mathbf{s}_{i}^{(1)}$.
+We leave for future work the comparison of the approximated analytical solution of Eq.
+(16) with the fully numerical solution of Eq.
+(C.2).
+IV Results
+IV.1 Dark ages
+During the dark ages, the only relevant matter species that contributes to the 21-cm signal are the baryons, as they source the emission and absorption of 21-cm radiation by hydrogen atoms, the recombination rate that governs $T_{k}$, and the coupling of $T_{s}$ to $T_{k}$ via collisional excitations.
+Thus the fluctuations in the 21-cm signal during the dark ages are entirely proportional to $\delta_{b}$.
+Because $\delta_{b}\ll 1$ during the dark ages, the scale-dependent evolution of the baryon density field barely affects the global 21-cm signal, i.e.
+the sky-averaged brightness temperature as a function of redshift, $\bar{T}_{21}\left(z\right)$.
+Indeed, for the cosmological and astrophysical parameters assumed in this work, we find that the SIGF approximation yields a maximum error of 0.4% in $\bar{T}_{21}$ at $z\simeq 14$, where the global signal reaches its minimum value.
+Although the scale-dependent evolution of $\delta_{b}$ does not affect the global 21-cm signal, it does affect its two-point statistics. Specifically, let us consider the 21-cm power spectrum,
+$\Delta_{21}^{2}\left(k,z\right)=\frac{k^{3}\bar{T}_{21}^{2}\left(z\right)P_{21}\left(k,z\right)}{2\pi^{2}},$ (29)
+where $P_{21}\left(k,z\right)$ is the angle-averaged Fourier transform of the two-point correlation function $\langle\delta_{21}\left(\mathbf{x},z\right)\delta_{21}\left(\mathbf{x}^{\prime},z\right)\rangle$, while $\delta_{21}$ is the local contrast in the brightness temperature, $\delta_{21}\left(\mathbf{x},z\right)\equiv T_{21}\left(\mathbf{x},z\right)/\bar{T}_{21}\left(z\right)-1$.
+In a previous work Flitter and Kovetz (2024b), we have shown that during the dark ages $\Delta_{21}^{2}\left(k,z\right)\propto\mathcal{T}_{b}^{2}\left(k,z\right)\ll\mathcal{T}_{c}^{2}\left(k,z\right)$.
+In Fig.
+3, we show $\Delta_{21}^{2}\left(k,z\right)$, as calculated by 21cmFirstCLASS after the incorporation of the SDGF in the code.
+As expected, $\Delta_{21}^{2}\left(k,z\right)$ is suppressed during the dark ages when $\delta_{b}$ is correctly evolved with the SDGF.
+Specifically, we find that the SIGF approximation yields a $\sim 40\%$ error in the 21-cm power spectrum at $z=100$ for $k\gtrsim 0.1\,\mathrm{Mpc}^{-1}$.
+IV.2 Cosmic dawn and reionization
+As was previously discussed in Sec.
+II, when the Universe enters the cosmic dawn epoch, the dependence of the 21-cm signal on the matter fields becomes more complicated as (1) non-linear fluctuations in $\delta_{b}$ become non-negligible and (2) both $\delta_{b}$ and $\delta_{m}$ need to be taken into account.
+We show in the top panels of Fig.
+4 the comparison between the more precise $\delta_{b}^{\mathrm{NL}}$, computed with the scale-dependent first order LPT, and the common calculation that evolves the density field with the SIGF.
+At $z=20$, we see that the error is of the order of a percent at the most over- and under-dense regions.
+This error is expected to increase at higher redshifts due to the departure of the SDGF from the SIGF, while at lower redshifts it vanishes as $\mathcal{D}_{b}\left(k,z\right)\to D\left(z\right)$, see Fig.
+1.
+The error in the evaluation of $\delta_{b}^{\mathrm{NL}}$ with the SIGF is enhanced when the 21-cm power spectrum is considered, due to the intricate interplay between non-linear fluctuations in different fields, e.g.
+$T_{k}$ and $J_{\alpha}$.
+Indeed, we see from comparing between the black and blue curves in Fig.
+5 that the error in $\Delta_{21}^{2}$ may reach to $\mathcal{O}\left(10\%\right)$ at $z\gtrsim 12$.
+At low redshifts ($z\lesssim 10$), the error due to this approximation goes to zero because at these redshift $\delta_{b}$ approaches more closely to $\delta_{c}$.
+The importance of evolving the baryon density field non-linearly is evident in the bottom panels of Fig.
+4; while linear perturbation theory seems to give a rough estimate for the distribution of baryonic matter on very large scales, it fails to capture the clustering of baryons at low redshifts and small scales, and it wrongly predicts a more homogeneous density field.
+When computing the 21-cm power spectrum, the imprecise calculation from linear perturbation theory yields an order unity error at $z\lesssim 8$ at $k=0.3\,\mathrm{Mpc}^{-1}$, as demonstrated in Fig.
+5 (this effect however is also manifest on other scales, as can be seen in Fig.
+7).
+Importantly, this is the regime where upcoming 21-cm interferometers are mostly sensitive (c.f.
+Fig.
+7).
+Finally, Fig.
+5 also demonstrates how $\Delta_{21}^{2}\left(k,z\right)$ is modified due to the choice of working with either $\delta_{m}^{\mathrm{lin}}$ or $\delta_{m}^{\mathrm{NL}}$ in Eq.
+(II.4).
+Evidently, because the fluctuations in the brightness temperature are strongly connected to the fluctuations in the SFRD, this choice has large implications on the resulting 21-cm power spectrum during cosmic dawn.
+Figure 5: The 21-cm power spectrum as a function of redshift at $k=0.3\,\mathrm{Mpc}^{-1}$, when the baryon and the total matter over-density fields are evolved either linearly or non-linearly.
+The black curve corresponds to adopting the SIGF approximation in which $\delta_{b}=\delta_{c}=\delta_{m}$.
+The colors of the curves here match the same colors in Figs.
+6-7, namely the blue, orange and green curves can be associated with 21cmFAST, Zeus21-like, and 21cmFirstCLASS, respectively.
+Here though, unlike Figs.
+6-7, in all colorful curves we distinguish between baryons and CDM (i.e.
+in all colorful curves we use the SDGF), and we have the same $v_{cb}$ delaying effect (see more information at Sec.
+III).
+Figure 6: The global 21-cm signal under different assumptions.
+Green lines are obtained using 21cmFirstCLASS and separate the baryons and dark matter evolution; other lines assume $\delta_{b}\equiv\delta_{c}$.
+In the blue curve, we ran 21cmFAST under the default settings of non-linear evolution; the orange curve instead is obtained setting the density field to evolve linearly, similarly as in Zeus21 (the two codes well agree in these settings).
+Solid (dashed) lines correspond to codes that adopt the non-linear (linear) matter density field in their calculations.
+In both green curves of 21cmFirstCLASS, the baryon density field is evolved non-linearly.
+Figure 7: Comparison of the 21-cm power spectrum on different scales, under the same conditions as in Fig.
+6.
+Here, we show with error bars the expected noise from HERA in its design sensitivity at low redshifts ($z<10$) assuming the moderate foreground scenario, as evaluated by the 21cmSense code Pober et al.
+(2013, 2014) (we used the same 21cmSense settings as detailed in Ref.
+Flitter and Kovetz (2024a)).
+We also show a separate thick error bar at $z=14$, corresponding to a sensitivity estimate for SKA Meriot and Semelin (2024).
+V Discussion
+The results shown in the previous section demonstrate that differences in the simulated 21-cm signal can indeed arise by simply using different matter field configurations in the simulation.
+We now examine if these differences could be significant in future detections during cosmic dawn and reionization epochs.
+We thus compare between 21cmFirstCLASS and two popular codes used in the literature, 21cmFAST and Zeus21.
+Similarly to 21cmFirstCLASS, 21cmFAST performs 2LPT in order to compute the non-linear fluctuations in the matter fields, however there is no distinction between baryons and CDM and the code assumes $\delta_{b}=\delta_{c}=\delta_{m}$.
+Zeus21 also does not make the distinction between baryons and CDM, but unlike 21cmFAST it assumes the linear matter density field $\delta_{m}^{\mathrm{lin}}$ throughout the calculation.
+A summary of the differences between the codes is given in Table  1.
+Feature 21cmFAST Zeus21 21cmFirstCLASS Diff. in $\Delta_{21}^{2}\left(k,z\right)$ Diff. in $\Delta_{21}^{2}\left(k,z\right)$
+(cosmic dawn and reionization) (dark ages)
+Non-linear evolution $\sim 20\%\left(z=15,k=0.3\,\mathrm{Mpc}^{-1}\right)$
+Scale dependent growth $\sim 10\%\left(z=15,k=0.3\,\mathrm{Mpc}^{-1}\right)$ $\sim 40\%\left(z=100,k=0.3\,\mathrm{Mpc}^{-1}\right)$
+Linear theory in EPS $\sim 20\%\left(z=7,k=0.1\,\mathrm{Mpc}^{-1}\right)$
+Table 1: Summary of features studied in this work, comparing between three public codes: 21cmFAST, Zeus21 and 21cmFirstCLASS.
+The leftmost column shows the three features discussed in this work.
+“Non-linear evolution” corresponds to perturbing the total matter density field non-linearly, $\delta=\delta^{\mathrm{NL}}$.
+“Scale dependent growth” corresponds to the relaxation of the SIGF approximation, $\delta_{b}\neq\delta_{c}$.
+“Linear theory in EPS” corresponds to having $\delta_{m}=\delta_{m}^{\mathrm{lin}}$ in Eq.
+(II.4) while removing the non-linear $v_{cb}$ term in the calculation of $\sigma\left(M_{h},z\right)$ (see more details in Sec.
+III).
+The middle columns display whether or not these features are present in the three codes.
+The two rightmost columns show the relative difference in the 21-cm power spectrum due to switching off the features.
+During the dark ages, only the second feature affects the 21-cm power spectrum, thus 21cmFirstCLASS is the only code among the three that can simulate that epoch.
+The redshifts of our quoted benchmarks correspond to the minima of the global signal at $z=15$ and $z=100$, or the maximum at $z=7$.
+Before reionization ($z\gtrsim 7$), the relative differences do not vary significantly with scale between $0.1\leq k\leq 0.6\,\mathrm{Mpc}^{-1}$.
+The comparison between the codes is shown in Figs.
+6-7, where 21cmFAST (Zeus21) is represented with blue (orange) curve.
+As 21cmFirstCLASS is able to use either $\delta_{m}^{\mathrm{NL}}$ or $\delta_{m}^{\mathrm{lin}}$ (while always using the non-linear $\delta_{b}^{\mathrm{NL}}$ in the simulation after the dark ages are over), these cases are represented by the solid and dashed green curves, respectively.
+We note that in order to compare apples to apples, we obtained the orange “Zeus21-like” curve by running 21cmFAST with linear perturbation theory, which was shown to agree well with the Zeus21 code in Refs.
+Muñoz (2023); Cruz et al.
+(2024b).
+In all three codes, the evaluated global signal is roughly the same, though minor differences can still be observed in Fig.
+6.
+When 21cmFirstCLASS runs with the $\delta_{m}^{\mathrm{lin}}$ configuration, the signal is slightly shifted to higher redshifts, as $\sigma\left(M_{h},z\right)$ is calculated directly from linear perturbation theory, Eq.
+(20), without the non-linear $v_{cb}$ modification that suppresses power at small scales (see more details in Sec.
+III).
+Furthermore, the difference in the emission signal during the reionization epoch can get up to 12% when comparing 21cmFirstCLASS with $\delta_{m}^{\mathrm{lin}}$ to 21cmFAST (or to 21cmFirstCLASS with $\delta_{m}^{\mathrm{NL}}$).
+This discrepancy increases to 20% when Zeus21 is compared to 21cmFAST.
+These differences arise due to the highly non-linear fluctuations in that epoch; the fluctuations in the signal are so large that they do not cancel out when the mean is considered.
+Next, in Fig.
+7 we compare $\Delta_{21}^{2}\left(k,z\right)$, as evaluated by the three codes.
+In this figure, the $z\lesssim 10$ error bars indicate the expected noise level from HERA in its design sensitivity under the conservative assumption of a moderate foreground scenario, evaluated with 21cmSense Pober et al.
+(2013, 2014), while the $z=14$ error bar illustrates the sensitivity expected for SKA, estimated by Ref.
+Meriot and Semelin (2024).
+On large scales ($k\lesssim 0.3\,\mathrm{Mpc}^{-1}$), the 21-cm power spectrum features the same differences in the global signal at low redshifts, and the type of the assumed matter field ($\delta_{m}^{\mathrm{NL}}$ or $\delta_{m}^{\mathrm{lin}}$) is crucial for accurately simulating the signal within the noise level of HERA.
+This conclusion also holds for small scales ($k\gtrsim 0.3\,\mathrm{Mpc}^{-1}$), though to a lesser extent.
+Interestingly, at intermediate scales of $k\sim 0.3\,\mathrm{Mpc}^{-1}$ the type of the assumed matter field less matters, suggesting that at these scales the signal mostly depends on $\delta_{b}$ rather than $\delta_{m}$, though we comment that this conclusion might change if different astrophysical parameters are considered.
+In all scales, the usage of $\delta_{b}^{\mathrm{lin}}$ is inappropriate for simulating the signal within HERA sensitivity range, especially during the epoch of reionization.
+At higher redshifts, we see that distinguishing $\delta_{b}$ from $\delta_{c}$ causes differences in $\Delta_{21}^{2}\left(k,z\right)$ on large scales that are comparable to the expected SKA noise level.
+Fig.
+7 thus demonstrates that the subtleties discussed in this work should matter in future analyses of 21-cm signal measurements.
+Yet, perhaps the most crucial subtlety in simulating the signal may come from another nuance which is treated the same in all three codes discussed here and has to do with the appearance of the smoothing radius $R$ in the RHS of Eq.
+(II.4).
 According to EPS, this $R$ should correspond to the region in which the local collapsed fraction is evaluated.
@@
 Note that the cell size cannot be arbitrarily small as we ought to have $M_{\mathrm{min}}<M_{R}$ for a non-zero local collapsed fraction (see square root in Eq.
-(III.2)).
+II.4).
 However, this condition is satisfied for a typical simulation cell of $R\sim L_{\mathrm{cell}}\sim 1\,\mathrm{Mpc}$; according to Eq.
@@
 […]
-V Conclusions
-Building a robust theoretical framework for the 21-cm signal is challenging, especially during the cosmic dawn and reionization epochs, due to the involved complicated astrophysics of galaxy evolution and star formation.
-While former studies of the 21-cm signal have focused on refining the astrophysical models, they largely glossed over a trivial but crucial aspect of the signal: it is sourced from neutral hydrogen atoms in the IGM between recombination and reionization, and these hydrogen atoms are baryons.
-Therefore, even though baryons are coupled gravitationally to CDM at low redshifts, the relevant density field to consider is not the CDM density field, but rather the baryon density field.
-Approximating the baryon density field with the CDM density field when tracing the evolution of the neutral hydrogen atoms in the IGM, leads to misestimation of the 21-cm signal.
-In this work, we distinguished between baryons and CDM to carefully study the cosmological impact of the baryons on the 21-cm signal.
+VI Conclusions
+Building a robust theoretical framework for the 21-cm signal is challenging, especially during the cosmic dawn and reionization epochs.
+Although valuable tools are available in the literature, all of them involve approximations or potential misinterpretations, often without thoroughly exploring their implications or consequences.
+For example, most of the codes model the 21-cm signal tracing the non-linear CDM density field, modeling its evolution using the SIGF.
+These choices glosses over a trivial but crucial aspect: since the 21-cm signal is sourced from neutral hydrogen atoms in the IGM, its modeling should follow the baryon density field.
+Even though baryons are coupled gravitationally to CDM at low redshifts, at higher redshifts this fact breaks down and the evolution of the two fields has to be carried on independently, also accounting for scale-dependent effects.
+Neglecting all of this potentially leads to misestimation of the 21-cm signal.
+In this work, we compared two widely used public 21-cm codes, 21cmFAST Mesinger et al.
+(2011); Muñoz et al.
+(2022) and Zeus21 Muñoz (2023); Cruz et al.
+(2024b), with our code, 21cmFirstCLASS Flitter and Kovetz (2024a, b), an extension to 21cmFAST that enables to begin the simulation from recombination with any desired cosmology.
+We outlined the main approximations introduced by the former and relaxed them in our implementation, demonstrating their impact on the 21-cm global signal and power spectrum.
+We distinguished between baryons and CDM to carefully study the cosmological impact of the baryons on the 21-cm signal, similarly as in Ref.
+Naoz and Barkana (2005), but we also extended our analysis to lower redshifts, after galaxies were formed.
 Unlike most studies which have adopted the SIGF approximation, whereby $\delta_{b}\propto D\left(z\right)$ and evolves in a scale-independent manner, we computed the baryon density field more precisely in 21cmFirstCLASS using the SDGF $\mathcal{D}_{b}\left(k,z\right)$.
@@
 We then moved to the non-linear epochs of cosmic dawn and reionization.
-While the exact physics in these epochs is still under uncertainty, especially as relates to the efficiency and production rate of stars and galaxies, two facts are certain regarding the cosmological role of the baryons in shaping the 21-cm signal: (1) their relative velocity with respect to CDM delays the signal, and (2) their non-linear clustering cannot be neglected at low redshifts and on small scales.
+As expected, even though some models of the SFRD consider the “linear” $\delta_{m}^{\mathrm{lin}}$ as an input to the conditional HMF, it is inconsistent to consider the “linear” $\delta_{b}^{\mathrm{lin}}$ once the dark ages are over, as this yields errors that are beyond the noise-level of 21-cm interferometers such as HERA.
+This is particularly true at low redshifts ($z\lesssim 8$) and large scales ($k\lesssim 0.3\,\mathrm{Mpc}^{-1}$), where the sensitivity of the interferometers is at maximum.
+Furthermore, we find that the SIGF approximation works very well for $z\lesssim 12$, as this is the regime where the baryons and CDM can indeed be characterized as a single fluid.
+In contrast, at higher redshifts, the SIGF approximation may lead to $\mathcal{O}\left(10\%\right)$ error, implying that advanced detectors like SKA will be sensitive enough such that the SIGF approximation is inadequate for simulating the signal in the required precision, even during cosmic dawn.
 In addition to understanding how the 21-cm signal depends on the baryon over-density $\delta_{b}$, we also discussed the role of the total matter over-density $\delta_{m}$ in the calculation.
 This enters through the collapsed fraction, Eq.
-(III.2), as part of the EPS formalism.
+(II.4), as part of the EPS formalism.
 While it is clear that the 21-cm signal at cosmic dawn and reionization should depend on the “non-linear” $\delta_{b}$ (i.e., $\delta_{b}$ as evaluated by non-linear theory), two common approaches are considered in the literature for the total matter density field $\delta_{m}$.
@@
 While Zeus21 (21cmFAST) adopts the former (latter) approach, 21cmFirtCLASS enables the user to decide which $\delta_{m}$ should be used in Eq.
-(III.2).
+(II.4).
 Regardless, as was discussed in the previous section, all codes still make in their calculations some assumptions whose validity should be further examined in the future.
-Efforts to measure the cosmological 21-cm signal are being carried out by various experiments.
-The expected data from these experiments, together with new photometric images from the James Webb Space Telescope Gardner et al.
-(2006); Mirocha and Furlanetto (2019); Harikane et al.
-(2023); Naidu et al.
-(2022); Castellano et al.
-(2022); Adams et al.
-(2023); Yan et al.
-(2023); Hassan et al.
-(2023), present an exciting opportunity to probe uncharted regions of our Universe, thereby deepening our understanding of the physical processes that governed its early evolution.
-Such an ambitious goal cannot be fulfilled without having robust theoretical predictions to be matched with the data.
-We thus believe that this work may serve to bring us one step closer towards this goal.
+In conclusion, we can summarize our detailed comparison between the three codes as follows (see also Fig.
+7 and Table 1).
+Zeus21 is a speedy code that attempts to reproduce the same 21-cm power spectrum as in 21cmFAST (albeit on much shorter timescales), but it only works with linear theory for the density field.
+While it is debatable whether or not it is more appropriate to consider the linear matter density field $\delta_{m}^{\mathrm{lin}}$ in the EPS and SFRD calculations, it is clear that it is inconsistent to consider the linear baryon density field $\delta_{b}^{\mathrm{lin}}$ at low redshifts, where HERA and other 21-cm interferometers are most sensitive.
+Even though the exact separation between baryons and CDM is not critical for analyzing the data from current interferometers, it would become increasingly more important in the future when data from next generation interferometers like SKA becomes available.
+At the moment, 21cmFirstCLASS is the only semi-numerical code that treats separately and consistently the different evolution in the baryon and CDM density fields during cosmic dawn and reionization epochs.
+Since 21cmFirstCLASS is also capable of simulating the dark ages, where this effect is most pronounced, it is thus poised to serve as an important tool in analyzing the 21-cm signal as measured by future interferometers, whether they are ground-based or lunar-based.
 Acknowledgements.
 It is our pleasure to thank Julian B.
-Muñoz for reviewing the manuscript and sharing his comments.
+Muñoz for reviewing the manuscript and making comments that improved the quality of this paper.
+We wish to thank the anonymous referee for providing a thorough and insightful report.
 We would also like to thank Andrei Mesinger and Tal Adi for useful discussions.
@@
 3156/23).
-Appendix A Scale-dependent matter evolution
-Even though CDM is considered as a collisionless fluid in $\Lambda$CDM, the evolution of its density fluctuations is not perfectly scale invariant, due to the gravitational coupling with baryons.
-Indeed, at low redshifts ($z\lesssim 1$) $\delta_{c}$ and $\delta_{b}$ approach each other, and BAO features from $\delta_{b}$ can be seen in the CDM transfer function $\mathcal{T}_{c}\left(k,z\right)$.
-This is no longer true at the high redshifts relevant for the 21-cm signal.
-Thus, the commonly used linear approximation of $\delta_{c}\left(\mathbf{k},z\right)\approx D\left(z\right)\delta_{0}\left(\mathbf{k}\right)$ results in fake BAO features in $\delta_{c}\left(\mathbf{k},z\right)$ at high redshifts ($z\gtrsim 1$).
-This can be remedied by using the appropriate SDGF for CDM $\delta_{c}\left(\mathbf{k},z\right)\approx\mathcal{D}_{c}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where $\mathcal{D}_{c}\left(k,z\right)$ is defined in Eq.
-(10).
-We compare between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ in Fig.
-10.
-From the left panel, it indeed appears as if $\mathcal{D}_{c}\left(k,z\right)$ is scale-invariant, unlike $\mathcal{D}_{b}\left(k,z\right)$ (c.f.
-Fig.
-3).
-However, a closer comparison between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ on the right panel shows that the two can differ by more than 2% at high redshifts ($z\gtrsim 25$), especially at small scales ($k\gtrsim 1\,\mathrm{Mpc}^{-1}$).
-Moreover, the fake BAO that the SIGF approximation yields is clearly seen on the right panel, through the non-monotonous behavior of $\mathcal{D}_{c}\left(k,z\right)/D\left(z\right)$ when $k$ is increased.
-Figure 10: Left panel: the scale-dependent growth factor for CDM.
-Right panel: comparison with the scale-independent growth factor.
-The oscillations with scale on the right panel indicate fake BAOs feature when the SIGF approximation is used.
-Figure 11: Top panels: The ratio of $\sigma\left(M_{h},z\right)$ and the Sheth-Tormen HMF, when they are calculated with the SIGF and the SDGF.
-Bottom left panel: The ratio of the collapsed fraction of halos that host ACGs (with $M_{\mathrm{min}}$ given by Eq.
-(IV)), when calculated with the SIGF and the SDGF.
-Bottom right panel: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the matter over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
-As we discussed in Sec.
-III.2, the local SFRD can be computed with the EPS formalism, where the local linear matter over-density $\delta_{m}^{\mathrm{lin}}\left(\mathbf{x},z\right)$ is considered.
-This quantity can be computed in Fourier space via $\delta^{\mathrm{lin}}_{m}\left(\mathbf{k},z\right)\approx\mathcal{D}_{m}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where
-$\mathcal{D}_{m}\left(k,z\right)\equiv\frac{\mathcal{T}_{m}\left(k,z\right)}{\mathcal{T}_{m}\left(k,z=0\right)}\approx\frac{\Omega_{b}}{\Omega_{m}}\frac{\mathcal{T}_{b}\left(k,z\right)}{\mathcal{T}_{b}\left(k,z=0\right)}+\frac{\Omega_{c}}{\Omega_{m}}\frac{\mathcal{T}_{c}\left(k,z\right)}{\mathcal{T}_{c}\left(k,z=0\right)}=\frac{\Omega_{b}}{\Omega_{m}}\mathcal{D}_{b}\left(k,z\right)+\frac{\Omega_{c}}{\Omega_{m}}\mathcal{D}_{c}\left(k,z\right),$ (25)
-where we used the excellent approximation that $\mathcal{T}_{b}$, $\mathcal{T}_{c}$ and $\mathcal{T}_{m}$ are all the same at $z=0$.
-This definition allows us to write $\sigma^{2}\left(M_{h},z\right)$ from Eq.
-(19) as
-$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z=0\right)\mathcal{D}_{m}^{2}\left(k,z\right)W_{\mathrm{ball}}^{2}\left(kR_{M_{h}}\right).$ (26)
-In the SIGF approximation, it is assumed that $\mathcal{D}_{m}\left(k,z\right)\approx D\left(z\right)$ and hence $\sigma\left(M_{h},z\right)\propto D\left(z\right)$.
-Therefore, the fake BAOs induced by the SIGF approximation at high redshifts cause an error of a few percent in $\sigma\left(M_{h},z\right)$, mainly in low halo masses, and an order unity error in the derived HMF, as we show in the top panels of Fig.
-11.
-In making the comparison for the HMF though, it is important to bare in mind that the ST parameters were fitted to the HMF that is computed with the SIGF approximation.
-It is thus intriguing to wonder if the output from N-body simulations can be better fitted by relaxing the SIGF approximation.
-When the collapsed fraction in ACGs is considered, the lower limit of the integral in Eq.
-(20) is at $M_{h}\sim 10^{8}\,M_{\odot}$, which is coincidently the regime where the error in the HMF is minimum, while higher halo masses less contribute to the integral.
-Thus, the error in the collapsed fraction of halos that host ACGs is sub-percent, as shown in the bottom left panel in Fig.
-11.
-When MCGs are considered too, their $M_{\mathrm{min}}$ is smaller than the equivalent quantity in ACGs by an order of magnitude, and it is subject to inhomogeneous baryonic feedback such as LW radiation and relative velocity with respect to CDM.
-We compare in the last panel of Fig.
-11 how the SIGF approximation, applied both in ACGs and MCGs, affects the 21-cm power spectrum.
-As it appears, the error introduced by the SIGF approximation in the HMF at high redshifts has almost no effect on the 21-cm power spectrum.
-This conclusion however relies on the assumed astrophysical model and the fiducial astrophysical parameters we have chosen.
-In other models, where the star production rate is predicted to be more efficient at high redshifts, the SIGF approximation may yield larger errors in the 21-cm signal.
-We defer this investigation for future work.
-Appendix B Boltzmann equation for the spin temperature
+Appendix A Boltzmann equation for the spin temperature
 Eq.
@@
 $\frac{d}{dt}\left(\frac{n_{1}}{n_{0}}\right)=-3T_{*}\frac{dT_{s}^{-1}}{dt}.$ (41)
-Combining Eqs. (B)-(37) thus yields
+Combining Eqs. (A)-(41) thus yields
 $\frac{dT_{s}^{-1}}{dt}=4\left(T_{k}^{-1}-T_{s}^{-1}\right)C_{10}+4\left(T_{\alpha}^{-1}-T_{s}^{-1}\right)P_{10}+\frac{4}{T_{*}}\left(1-\frac{T_{\gamma}}{T_{s}}\zeta_{\mathrm{CMB}}\right)A_{10}.$ (42)
@@
 We plot $\epsilon_{s}\left(z\right)$ during the dark ages in Fig.
-12.
+8.
 We find that this quantity reaches maximum at $\epsilon_{s,\max}\left(z=69\right)\approx 10^{-3}$.
@@
 (13), is valid also during the dark ages with excellent precision.
-Appendix C Temperature fluctuations in the dark ages
-In Ref.
-Flitter and Kovetz (2024b) we discussed the effect of kinetic temperature fluctuations on the brightness temperature fluctuations during the dark ages.
-Using linear perturbation theory, we showed that the two kinds of fluctuations are approximately related in Fourier space via
-$\delta_{21,\mathrm{iso}}\left(k,z\right)\approx\frac{2\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\delta_{b}\left(k,z\right)+\left[\frac{\bar{x}_{\mathrm{CMB}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\frac{\partial\ln\bar{\kappa}_{1-0}^{\mathrm{HH}}}{\partial\ln\bar{T}_{k}}-\frac{1}{1-\bar{T}_{k}/T_{\gamma}}\right]\delta_{T}\left(k,z\right),$ (43)
-where $\kappa_{1-0}^{\mathrm{HH}}$ is the collision rate (in units of $\mathrm{cm^{3}/sec}$) of hydrogen atoms with themselves.
-Here, $\delta_{T}\left(k,z\right)$ and $\delta_{21,\mathrm{iso}}\left(k,z\right)$ are the Fourier transform of the fractional fluctuation in $T_{k}\left(\mathbf{x},z\right)$ and $T_{21}\left(\mathbf{x},z\right)$, respectively, and the latter does not contain fluctuations sourced from peculiar velocity (see Eq.
-(12)).
-At the dark ages, the fluctuations $\delta_{T}$ are sourced by the perturbations in the baryon density field $\delta_{b}$, and thus it is useful to define the so called “adiabatic index”,
-$c_{T}\left(k,z\right)\equiv\frac{\delta_{T}\left(k,z\right)}{\delta_{b}\left(k,z\right)}.$ (44)
-The reason for this name is because for small $\delta_{b}\left(k,z\right)$ we can write $1+\delta_{T}\left(k,z\right)=\left[1+\delta_{b}\left(k,z\right)\right]^{c_{T}\left(k,z\right)}$.
-This adiabatic index can be calculated analytically by solving simultaneously the linearly perturbed differential equations for $T_{k}$ and $x_{e}$.
-We show in Fig.
-13 the analytical solution for $c_{T}\left(k,z\right)$ as a function of redshift for $k=0.5\,\mathrm{Mpc}^{-1}$.
-The blue dashed curve displays the scale-independent $c_{T}\left(z\right)$ which is obtained under the very crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, namely $\delta_{b}\left(k,z\right)$ evolves in a scale-independent manner.
-In the solid blue curve we have relaxed this assumption and used the exact $\mathcal{D}_{b}\left(k,z\right)$, as defined in Eq.
+Appendix B Scale-dependent matter evolution
+Even though CDM is considered as a collisionless fluid in $\Lambda$CDM, the evolution of its density fluctuations is not perfectly scale invariant, due to the gravitational coupling with baryons.
+Indeed, at low redshifts ($z\lesssim 1$) $\delta_{c}$ and $\delta_{b}$ approach each other, and BAO features from $\delta_{b}$ can be seen in the CDM transfer function $\mathcal{T}_{c}\left(k,z\right)$.
+This is no longer true at the high redshifts relevant for the 21-cm signal.
+Thus, the commonly used linear approximation of $\delta_{c}\left(\mathbf{k},z\right)\approx D\left(z\right)\delta_{0}\left(\mathbf{k}\right)$ results in fake BAO features in $\delta_{c}\left(\mathbf{k},z\right)$ at high redshifts ($z\gtrsim 1$).
+This can be remedied by using the appropriate SDGF for CDM $\delta_{c}\left(\mathbf{k},z\right)\approx\mathcal{D}_{c}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where $\mathcal{D}_{c}\left(k,z\right)$ is defined in Eq.
 (14).
-For both curves, we assumed a recombination rate from the RECFAST code Seager et al.
-(1999, 2000).
-Evidently, the scale-dependent $c_{T}\left(k,z\right)$ curve reaches higher values than the scale-independent curve.
-To verify the consistency of the calculations of 21cmFirstCLASS at the dark ages we also show in Fig.
-13 the output of the code for $c_{T}\left(k=0.5\,\mathrm{Mpc}^{-1},z\right)$.
-This output was achieved by computing the $c_{T}\left(\mathbf{k},z_{i}\right)$ coeval box (in Fourier space) at every redshift iteration $z_{i}$.
-Because of the scale-dependence, this box is not homogeneous, and so we need to select the appropriate cells that correspond to a desired wavenumber $k_{0}$ (in Fig.
-13, $k_{0}=0.5\,\mathrm{Mpc}^{-1}$).
-We do this selection process using the following criterion,
-$k_{0}-\Delta k/2<|\mathbf{k}|<k_{0}+\Delta k/2,$ (45)
-where $\Delta k=2\pi/L_{\mathrm{box}}$ and $L_{\mathrm{box}}$ is the box size of the simulation.
-Once we have the appropriate cells from the $c_{T}\left(\mathbf{k},z_{i}\right)$ box, we take their median, depicted by the orange curve in Fig.
-13.
-There is an excellent agreement between the analytical calculation and the numerical calculation from 21cmFirstCLASS, even though in the former we have assumed the recombination rate from RECFAST, while the latter works with HYREC Ali-Haimoud and Hirata (2011); Lee and Ali-Haïmoud (2020).
-Figure 13: The scale-dependent adiabatic index for $k=0.5\,\mathrm{Mpc}^{-1}$ as a function of redshift during the dark ages.
-The blue curves show the analytic calculation, based on the equations found in Ref.
-Flitter and Kovetz (2024b), while the orange curve shows the numerical output from 21cmFirstCLASS, see text for details.
-For the dashed blue curve, we have taken the crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, yielding an inconsistent adiabatic index.
-Appendix D Scale-dependent Lagrangian perturbation theory
+We compare between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ in Fig.
+9.
+From the left panel, it indeed appears as if $\mathcal{D}_{c}\left(k,z\right)$ is scale-invariant, unlike $\mathcal{D}_{b}\left(k,z\right)$ (c.f.
+Fig.
+1).
+However, a closer comparison between $\mathcal{D}_{c}\left(k,z\right)$ and $D\left(z\right)$ on the right panel shows that the two can differ by more than 2% at high redshifts ($z\gtrsim 25$), especially at small scales ($k\gtrsim 1\,\mathrm{Mpc}^{-1}$).
+Moreover, the fake BAO that the SIGF approximation yields is clearly seen on the right panel, through the non-monotonous behavior of $\mathcal{D}_{c}\left(k,z\right)/D\left(z\right)$ when $k$ is increased.
+As we discussed in Sec.
+II.5, the local SFRD can be computed with the EPS formalism, where the local linear matter over-density $\delta_{m}^{\mathrm{lin}}\left(\mathbf{x},z\right)$ is considered.
+This quantity can be computed in Fourier space via $\delta^{\mathrm{lin}}_{m}\left(\mathbf{k},z\right)\approx\mathcal{D}_{m}\left(k,z\right)\delta_{0}\left(\mathbf{k}\right)$, where
+$\mathcal{D}_{m}\left(k,z\right)\equiv\frac{\mathcal{T}_{m}\left(k,z\right)}{\mathcal{T}_{m}\left(k,z=0\right)}\approx\frac{\Omega_{b}}{\Omega_{m}}\frac{\mathcal{T}_{b}\left(k,z\right)}{\mathcal{T}_{b}\left(k,z=0\right)}+\frac{\Omega_{c}}{\Omega_{m}}\frac{\mathcal{T}_{c}\left(k,z\right)}{\mathcal{T}_{c}\left(k,z=0\right)}=\frac{\Omega_{b}}{\Omega_{m}}\mathcal{D}_{b}\left(k,z\right)+\frac{\Omega_{c}}{\Omega_{m}}\mathcal{D}_{c}\left(k,z\right),$ (47)
+where we used the excellent approximation that $\mathcal{T}_{b}$, $\mathcal{T}_{c}$ and $\mathcal{T}_{m}$ are all the same at $z=0$.
+This definition allows us to write $\sigma^{2}\left(M_{h},z\right)$ from Eq.
+(20) as
+$\sigma^{2}\left(M_{h},z\right)=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\star}}\right)^{n_{s}-1}\mathcal{T}^{2}_{m}\left(k,z=0\right)\mathcal{D}_{m}^{2}\left(k,z\right)W^{2}\left(kR_{M_{h}}\right).$ (48)
+Figure 9: Left panel: the scale-dependent growth factor for CDM.
+Right panel: comparison with the scale-independent growth factor.
+The oscillations with scale on the right panel indicate fake BAOs feature when the SIGF approximation is used.
+Figure 10: Top panels: The ratio of $\sigma\left(M_{h},z\right)$ and the Sheth-Tormen HMF, when they are calculated with the SIGF and the SDGF.
+Bottom left panel: The ratio of the collapsed fraction of halos that host ACGs (with $M_{\mathrm{min}}$ given by Eq.
+V), when calculated with the SIGF and the SDGF.
+Bottom right panel: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the matter over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
+In the SIGF approximation, it is assumed that $\mathcal{D}_{m}\left(k,z\right)\approx D\left(z\right)$ and hence $\sigma\left(M_{h},z\right)\propto D\left(z\right)$.
+Therefore, the fake BAOs induced by the SIGF approximation at high redshifts cause an error of a few percent in $\sigma\left(M_{h},z\right)$, mainly in low halo masses, and an order unity error in the derived HMF, as we show in the top panels of Fig.
+10.
+In making the comparison for the HMF though, it is important to bare in mind that the ST parameters were fitted to the HMF that is computed with the SIGF approximation.
+It is thus intriguing to wonder if the output from N-body simulations can be better fitted by relaxing the SIGF approximation.
+When the collapsed fraction in ACGs is considered, the lower limit of the integral in Eq.
+(21) is at $M_{h}\sim 10^{8}\,M_{\odot}$, which is coincidently the regime where the error in the HMF is minimum, while higher halo masses less contribute to the integral.
+Thus, the error in the collapsed fraction of halos that host ACGs is sub-percent, as shown in the bottom left panel in Fig.
+10.
+When MCGs are considered too, their $M_{\mathrm{min}}$ is smaller than the equivalent quantity in ACGs by an order of magnitude, and it is subject to inhomogeneous baryonic feedback such as LW radiation and relative velocity with respect to CDM.
+We compare in the last panel of Fig.
+10 how the SIGF approximation, applied both in ACGs and MCGs, affects the 21-cm power spectrum.
+As it appears, the error introduced by the SIGF approximation in the HMF at high redshifts has almost no effect on the 21-cm power spectrum.
+This conclusion however relies on the assumed astrophysical model and the fiducial astrophysical parameters we have chosen.
+In other models, where the star production rate is predicted to be more efficient at high redshifts, the SIGF approximation may yield larger errors in the 21-cm signal.
+We defer this investigation for future work.
+Appendix C Scale-dependent Lagrangian perturbation theory
 Normally, in Lagrangian perturbation theory Friedman and Schutz (1978); Buchert (1989); Moutarde et al.
@@
 Equations (56), (57), (59) form a set of non-linear equations for $S_{i}$. We use perturbative approach to solve them.
-D.1 First order - Zel’dovich approximation
+C.1 First order - Zel’dovich approximation
 Let us expand $S_{i}\left(\mathbf{q},z\right)$ to its first order, $S_{i}\left(\mathbf{q},z\right)\approx S_{i}^{(1)}\left(\mathbf{q},z\right)$. Then, to first order,
@@
 where we used the linear expression for $\delta_{i}$ (Eq.
-(9)) as we seek first order solution to $S_{i}$.
+13) as we seek first order solution to $S_{i}$.
 To see that this solution is consistent, we expand the Euler-Poisson equation, Eq.
@@
 Eq.
-(7)).
+11).
 Furthermore, the homogeneity assumption we made in Eq.
@@
 as commonly done in the literature Efstathiou et al. (1985); Sirko (2005); Mesinger and Furlanetto (2007).
-D.2 Second order - 2LPT
+C.2 Second order - 2LPT
 Next, we expand $S_{i}$ to its second order, $S_{i}\left(\mathbf{q},z\right)\approx S_{i}^{(1)}\left(\mathbf{q},z\right)+S_{i}^{(2)}\left(\mathbf{q},z\right)$.
@@
 If variable separation was possible, then from Eq.
-(D.2) we could have obtained two different equations, one for the spatial part of $S_{i}^{(2)}$, and one for its temporal part.
+(C.2) we could have obtained two different equations, one for the spatial part of $S_{i}^{(2)}$, and one for its temporal part.
 However, as we established in Eq.
@@
 Note that given $S_{i}^{(1)}$, Eq.
-(D.2) does not give a differential equation for $S_{i}^{(2)}$, but rather for the Laplacian of $S_{i}^{(2)}$.
+(C.2) does not give a differential equation for $S_{i}^{(2)}$, but rather for the Laplacian of $S_{i}^{(2)}$.
 In order to obtain an integro-differential equation for $S_{i}^{(2)}$, we take the Fourier transform of Eq.
-(D.2).
+(C.2).
 While applying Fourier transform on the LHS gives a multiplication by $-k^{2}$, taking the Fourier transform of the RHS, being proportional to the square of second derivatives of $S_{i}^{(1)}$, would result in a convolution.
@@
 This integro-differential equation actually contains two equations, one for the baryons ($i=b$) and one for CDM ($i=c$), and these two equations have to be simultaneously solved due to the gravitational coupling between the species, while imposing zero initial conditions for $S_{i}^{\left(2\right)}$ and its derivative at $z=z_{0}$.
-D.2.1 Scale independent growth
+C.2.1 Scale independent growth
 While in principle Eq.
-(D.2) can be solved numerically, this is computationally expensive, as for each cell $\mathbf{q}$ in the grid, double integration (that involves all the cells in the box in Fourier space) is required, for both species.
+(C.2) can be solved numerically, this is computationally expensive, as for each cell $\mathbf{q}$ in the grid, double integration (that involves all the cells in the box in Fourier space) is required, for both species.
 To ease the computational cost, we will make the approximation that is done vastly in the literature, of approximating $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$, i.e., we replace the SDGF for each species with a universal SIGF.
@@
 Plugging this universal solution back to Eq.
-(D.2) yields
+(C.2) yields
 $\ddot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)+2H\dot{S}_{i}^{\left(2\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}\sum_{j}\frac{\Omega_{j}}{\Omega_{m}}S_{j}^{\left(2\right)}\left(\mathbf{q},z\right)$
@@
 Because the RHS of Eq.
-(D.2.1) is universal for all species, and since $S_{i}^{(2)}$ has zero initial conditions at $z=z_{0}$ for every species $i$, we deduce that the SIGF approximation implies that $S_{i}^{(2)}\left(\mathbf{q},z\right)$ also has a universal solution $S^{(2)}\left(\mathbf{q},z\right)$.
+(C.2.1) is universal for all species, and since $S_{i}^{(2)}$ has zero initial conditions at $z=z_{0}$ for every species $i$, we deduce that the SIGF approximation implies that $S_{i}^{(2)}\left(\mathbf{q},z\right)$ also has a universal solution $S^{(2)}\left(\mathbf{q},z\right)$.
 This property can be further inferred for any higher order in perturbation theory.
 Hence, the SIGF approximation corresponds to solving the perturbed equations of motions under the approximation $\delta_{b}\approx\delta_{c}\approx\delta_{m}$, namely we directly solve for the total matter density field.
-After substituting the universal solution $S_{i}^{(2)}\left(\mathbf{q},z\right)=S^{(2)}\left(\mathbf{q},z\right)$, Eq. (D.2.1) can be written in the following form,
+After substituting the universal solution $S_{i}^{(2)}\left(\mathbf{q},z\right)=S^{(2)}\left(\mathbf{q},z\right)$, Eq. (C.2.1) can be written in the following form,
 $\nabla^{2}_{\mathbf{q}}\left[\ddot{S}^{\left(2\right)}\left(\mathbf{q},z\right)+2H\dot{S}^{\left(2\right)}\left(\mathbf{q},z\right)-4\pi G\bar{\rho}_{m}S^{\left(2\right)}\left(\mathbf{q},z\right)\right]=-4\pi G\bar{\rho}_{m}D^{2}\left(z\right)\left[\frac{1}{2}\left(\nabla_{\mathbf{q}}^{2}\psi^{\left(1\right)}\right)^{2}-\frac{1}{2}\left(\frac{\partial^{2}\psi^{\left(1\right)}}{\partial q_{n}\partial q_{m}}\right)^{2}\right].$ (73)
@@
 This closed solution for the displacement vector allows one to compute $\mathbf{s}^{\left(2\right)}$ for any $z$, independently from its history.
-D.2.2 Solution to $E\left(z\right)$
+C.2.2 Solution to $E\left(z\right)$
 Before solving for $E\left(z\right)$, we shall solve first the differential equation for the SIGF $D\left(z\right)$.
@@
 where the first equality holds to a percent level.
-Figure 14: Left panel: realization of the baryon over-density field at $z=20$, evolved non-linearly with scale-dependent 2LPT.
-Right panel: the fractional error in the normalized density field ($\Delta_{i}\equiv 1+\delta_{i}$) due to approximating $\delta_{b}\approx\delta_{c}$.
-Figure 15: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$ (solid curves), $k=0.3\,\mathrm{Mpc}^{-1}$ (dashed curves), and $k=0.15\,\mathrm{Mpc}^{-1}$ (dotted curves), when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve).
-For fair comparison, in all curves we used the non-linear $\delta_{m}^{\mathrm{NL}}$ in the EPS formalism, as in standard 21cmFAST.
-D.3 The effect of scale-dependent growth on the 21-cm signal at cosmic dawn
-In 21cmFAST, the SIGF approximation of $\mathcal{D}_{i}\left(k,z\right)\approx D\left(z\right)$ is taken, and Eqs.
-(62), (74) are implemented in the code for computing the first and second order expressions for the displacement vector field.
-In order to compute the non-linear baryon density field more precisely, we replaced Eq.
-(62) in 21cmFirstCLASS with Eq.
-(61), i.e.
-we compute the first order expression for $\mathbf{s}_{i}^{(1)}$ via the SDGF $\mathcal{D}_{i}\left(\mathbf{k},z\right)$, for each species $i$.
-As for the second order expression, $\mathbf{s}_{i}^{(2)}$, we did not attempt to solve the coupled integro-differential equations given by Eq.
-(D.2), and instead we keep using the approximated Eq.
+Appendix D Temperature fluctuations in the dark ages
+In Ref.
+Flitter and Kovetz (2024b) we discussed the effect of kinetic temperature fluctuations on the brightness temperature fluctuations during the dark ages.
+Using linear perturbation theory, we showed that the two kinds of fluctuations are approximately related in Fourier space via
+$\delta_{21,\mathrm{iso}}\left(k,z\right)\approx\frac{2\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\delta_{b}\left(k,z\right)+\left[\frac{\bar{x}_{\mathrm{CMB}}}{\bar{x}_{\mathrm{CMB}}+\bar{x}_{\mathrm{coll}}}\frac{\partial\ln\bar{\kappa}_{1-0}^{\mathrm{HH}}}{\partial\ln\bar{T}_{k}}-\frac{1}{1-\bar{T}_{k}/T_{\gamma}}\right]\delta_{T}\left(k,z\right),$ (87)
+where $\kappa_{1-0}^{\mathrm{HH}}$ is the collision rate (in units of $\mathrm{cm^{3}/sec}$) of hydrogen atoms with themselves.
+Here, $\delta_{T}\left(k,z\right)$ and $\delta_{21,\mathrm{iso}}\left(k,z\right)$ are the Fourier transform of the fractional fluctuation in $T_{k}\left(\mathbf{x},z\right)$ and $T_{21}\left(\mathbf{x},z\right)$, respectively, and the latter does not contain fluctuations sourced from peculiar velocity (see Eq.
+2).
+At the dark ages, the fluctuations $\delta_{T}$ are sourced by the perturbations in the baryon density field $\delta_{b}$, and thus it is useful to define the so called “adiabatic index”,
+$c_{T}\left(k,z\right)\equiv\frac{\delta_{T}\left(k,z\right)}{\delta_{b}\left(k,z\right)}.$ (88)
+The reason for this name is because for small $\delta_{b}\left(k,z\right)$ we can write $1+\delta_{T}\left(k,z\right)=\left[1+\delta_{b}\left(k,z\right)\right]^{c_{T}\left(k,z\right)}$.
+This adiabatic index can be calculated analytically by solving simultaneously the linearly perturbed differential equations for $T_{k}$ and $x_{e}$.
+We show in Fig.
+11 the analytical solution for $c_{T}\left(k,z\right)$ as a function of redshift for $k=0.5\,\mathrm{Mpc}^{-1}$.
+The blue dashed curve displays the scale-independent $c_{T}\left(z\right)$ which is obtained under the very crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, namely $\delta_{b}\left(k,z\right)$ evolves in a scale-independent manner.
+In the solid blue curve we have relaxed this assumption and used the exact $\mathcal{D}_{b}\left(k,z\right)$, as defined in Eq.
 (14).
-This is justified as by construction the second order correction $\mathbf{s}_{i}^{(2)}$ has a smaller contribution to $\mathbf{s}_{i}$, compared to the first order correction $\mathbf{s}_{i}^{(1)}$.
-We leave for future work the comparison of the approximated analytical solution of Eq.
-(74) with the fully numerical solution of Eq.
-(D.2).
-We show in Fig.
-14 the comparison between the more precise $\delta_{b}^{\mathrm{NL}}$, computed with the scale-dependent first order LPT, and the common calculation that evolves the density field with the SIGF.
-At $z=20$, we see that the error is of the order of a percent at the most over- and under-dense regions.
-At higher redshifts, this error is expected to increase at higher redshifts due to the departure of the SDGF from the SIGF, while at lower redshifts it vanishes as $\mathcal{D}_{b}\left(k,z\right)\to D\left(z\right)$, see Fig.
-3.
-The error in the evaluation of $\delta_{b}^{\mathrm{NL}}$ with the SIGF is enhanced when the 21-cm power spectrum is considered, due to the intricate interplay between non-linear fluctuations in different fields, e.g.
-$T_{k}$ and $J_{\alpha}$.
-Indeed, we see in Fig.
-15 that the error in $\Delta_{21}^{2}$ may reach to $\mathcal{O}\left(10\%\right)$ at $z\gtrsim 12$.
-At low redshifts ($z\lesssim 10$), the error due to this approximation goes to zero because at these redshift $\delta_{b}$ approaches more closely to $\delta_{c}$.
+For both curves, we assumed a recombination rate from the RECFAST code Seager et al.
+(1999, 2000).
+Evidently, the scale-dependent $c_{T}\left(k,z\right)$ curve reaches higher values than the scale-independent curve.
+To verify the consistency of the calculations of 21cmFirstCLASS at the dark ages we also show in Fig.
+11 the output of the code for $c_{T}\left(k=0.5\,\mathrm{Mpc}^{-1},z\right)$.
+This output was achieved by computing the $c_{T}\left(\mathbf{k},z_{i}\right)$ coeval box (in Fourier space) at every redshift iteration $z_{i}$.
+Because of the scale-dependence, this box is not homogeneous, and so we need to select the appropriate cells that correspond to a desired wavenumber $k_{0}$ (in Fig.
+11, $k_{0}=0.5\,\mathrm{Mpc}^{-1}$).
+We do this selection process using the following criterion,
+$k_{0}-\Delta k/2<|\mathbf{k}|<k_{0}+\Delta k/2,$ (89)
+where $\Delta k=2\pi/L_{\mathrm{box}}$ and $L_{\mathrm{box}}$ is the box size of the simulation.
+Once we have the appropriate cells from the $c_{T}\left(\mathbf{k},z_{i}\right)$ box, we take their median, depicted by the orange curve in Fig.
+11.
+There is an excellent agreement between the analytical calculation and the numerical calculation from 21cmFirstCLASS, even though in the former we have assumed the recombination rate from RECFAST, while the latter works with HYREC Ali-Haimoud and Hirata (2011); Lee and Ali-Haïmoud (2020).
+Figure 11: The scale-dependent adiabatic index for $k=0.5\,\mathrm{Mpc}^{-1}$ as a function of redshift during the dark ages.
+The blue curves show the analytic calculation, based on the equations found in Ref.
+Flitter and Kovetz (2024b), while the orange curve shows the numerical output from 21cmFirstCLASS, see text for details.
+For the dashed blue curve, we have taken the crude approximation $\mathcal{D}_{b}\left(k,z\right)\approx D\left(z\right)$, yielding an inconsistent adiabatic index.